### v1

# Eigenvalue degeneracy in sparse random matrices

Abstract
In random matrices with independent and continuous matrix entries, the degeneracy probability of the eigenvalues is known to be zero. In this paper, random matrices including discontinuous matrix entries are analyzed in order to observe how degeneracy is generated. Using Erdös-Rényi matching probability theory of random bipartite graphs, we asymptotically evaluate the degeneracy probability of such random matrices. As a result, due to accumulation of the eigenvalues to the origin, a positive degeneracy probability is found for eigenvalues of a sparse random matrix model.
Graduate School of Mathematics, Nagoya University, Chikusa-ku,
Nagoya 464-8602, Japan
Keywords:
eigenvalue degeneracy, random matrices, sparse matrices, random graphs
MSC Classification: 60B20, 05C80
1 Introduction
In random matrix theory, eigenvalue distributions are often analyzed under the assumption that there is no eigenvalue degeneracy. This assumption enables diagonalizations of matrices with one-to-one correspondence of parameters, and the implicit function theorem is applied to observe the eigenvalue motion in reaction to a matrix motion. Such an argument is possible because there is a common sense that eigenvalue degeneracy of random matrices does not take place in almost all situations. As explained in Tao’s book[1] on random matrices, this common sense is based on the following fact.
Theorem 1.1
Let us set $N\geq 2$. In the space of $N\times N$ hermitian matrices, the codimension of the space of matrices with at least one eigenvalue degenerated is $3$. In the space of $N\times N$ real symmetric matrices, the codimension of the space of matrices with at least one eigenvalue degenerated is $2$.
This theorem appeared in the paper[2] by von Neumann and Wigner published in 1929. Let us explain it for a hermitian matrix. Suppose that an $N\times N$ hermitian matrix has $f$ distinct eigenvalues and the multiplicity of each distinct eigenvalue is $g_{1},\cdots,g_{f}$. Then such a hermitian matrix can be written in terms of $N^{2}+f-g_{1}^{2}-g_{2}^{2}-\cdots-g_{f}^{2}$ parameters. In a special case $g_{1}=g_{2}=\cdots=g_{f-1}=1$ with $g_{f}=2$, we need $N^{2}-3$ parameters. As a hermitian matrix in general is written in terms of $N^{2}$ parameters, the codimension is $3$. Therefore, if the probability measure of the whole space of hermitian matrices is absolutely continuous with respect to the $N^{2}$ dimensional Lebesgue measure, the degeneracy probability of the matrix eigenvalues is zero.
On the other hand, when the probability measure is discontinuous, the degeneracy probability of random matrix eigenvalues has been examined relatively recently. A paper[3] by Tao and Vu in 2017 considered the case when matrix entries (upper-triangular and diagonal entries) of a real symmetric matrix were independently distributed. The probability measure of each entry was not necessarily continuous. If the measure of a one point set was less than a constant independent of the matrix dimension $N$, the degeneracy probability of the eigenvalues converged to zero in the limit $N\rightarrow\infty$. Tao and Vu also studied Erdös-Rényi random graphs with $N$ vertices. When the probability to find a couple of vertices connected by an edge was independent of $N$, they showed that the eigenvalue degeneracy probability of the adjacency matrices approached to zero in the limit $N\rightarrow\infty$. Moreover, in the paper[4] by Luh and Vu in 2020, sparse $N\times N$ real symmetric matrices were investigated. In the limit $N\rightarrow\infty$, they assumed that the probability to have a zero matrix entry converged to one, and found a sufficient condition to make the eigenvalue degeneracy probability tend to zero.
In this paper, we analyze random matrix models with discontinuous probability measures in order to observe how degeneracy is generated. In particular, for a sparse random matrix model, we are able to find a positive degeneracy probability for the eigenvalues. Degeneracy occurs because of accumulation of the eigenvalues to the origin, and the accumulation comes from the discontinuity.
The outline of this paper is the following. In §2, we consider an $N\times N$ matrix which is an analytic function $\varphi(X_{1},X_{2},\cdots,X_{N})$ of $d$ independent and continuous random variables $X_{1},X_{2},\cdots,X_{d}$. When there exists at least one family of constants $x_{1},x_{2},\cdots,x_{d}$ making the $N$ eigenvalues of $\varphi(x_{1},x_{2},\cdots,x_{d})$ all distinct, then the $N$ eigenvalues of $\varphi(X_{1},X_{2},\cdots,X_{d})$ are all distinct with probability one. Thus we find that the eigenvalue degeneracy probability turns out to be either zero or one. This result can be applied to a variety of typical random matrix models including random unitary matrices.
In §3, matrix models related to bipartite graphs are introduced. In these models, if the $j$-th and $\ell$-th vertices are not connected by an edge in the bipartite graph, then the $(j,\ell)$ entry of the corresponding matrix is fixed to zero. The concept of perfect matching is borrowed from the theory of bipartite graphs and used in the study of degeneracy. If a bipartite graph has a perfect matching, the eigenvalues of at least one of the corresponding matrices do not degenerate. Such an analysis using the concept of graph theory is also applied to symmetric matrices.
In §4, employing the results of §2 and §3 together, we study the eigenvalue degeneracy probability for $N\times N$ sparse random matrices. The matrix entries are supposed to be continuously distributed expect at the origin. In order to evaluate the degeneracy probability, we refer to a theory by Erdös and Rényi concerning the perfect matching probability of random bipartite graphs. As a result, the degeneracy probability is asymptotically evaluated in the limit $N\rightarrow\infty$, and a positive degeneracy probability is worked out. Accumulation of the eigenvalues to the origin, due to the discontinuity of the matrix entry distribution, causes the degeneracy. In §5, we consider the degeneracy probability for symmetric sparse random matrices, and again find a positive asymptotic limit.
2 Eigenvalue degeneracy probability for continuous distributions
Matrix eigenvalues are roots of the characteristic equation. In order to study the root degeneracy, we first introduce the concept of a discriminant. Using discriminants, we show a sufficient condition for the eigenvalue degeneracy probability of random matrices to be zero or one.
Definition 2.1
For an $n$-th order polynomial
$p(\lambda)=\lambda^{n}+a_{n-1}\lambda^{n-1}+\cdots+a_{1}\lambda+a_{0},$ (2.1)
we set $b_{k}=ka_{k}$, $k=1,2,\cdots,n-1$. Then we define the discriminant $\Delta[p(\lambda)]$ of $p(\lambda)$ as the following determinant (multiplied with a coefficient)
$\Delta[p(\lambda)]=(-1)^{n(n-1)/2}\left|\begin{array}[]{ccccccccc}1&a_{n-1}&\cdots&\cdots&a_{1}&a_{0}&&&\\ &1&a_{n-1}&\cdots&\cdots&a_{1}&a_{0}&&\\ &&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\\ &&&1&a_{n-1}&\cdots&\cdots&a_{1}&a_{0}\\ n&b_{n-1}&\cdots&b_{2}&b_{1}&&&&\\ &n&b_{n-1}&\cdots&b_{2}&b_{1}&&&\\ &&\ddots&\ddots&\ddots&\ddots&\ddots&&\\ &&&n&b_{n-1}&\cdots&b_{2}&b_{1}&\\ &&&&n&b_{n-1}&\cdots&b_{2}&b_{1}\end{array}\right|$ (2.2)
of a $(2n-1)\times(2n-1)$ matrix. The upper n1n-1 rows consist of the coefficients $1,a_{n-1},a_{n-2},\cdots,a_{1},a_{0}$ of $p(\lambda)$ with one column shifted to the right each row. The lower $n$ rows consist of the coefficients $n,b_{n-1},b_{n-2},\cdots,b_{2},b_{1}$ of $p^{\prime}(\lambda)$ with one column shifted to the right each row. All vacancies are occupied with zeros.
Lemma 2.2
Let us suppose that the roots of an $n$-th order polynomial $p(\lambda)=\lambda^{n}+a_{n-1}\lambda^{n-1}+\cdots+a_{1}\lambda+a_{0}$ are $\lambda_{1},\lambda_{2},\cdots,\lambda_{n}$. Then the discriminant $\Delta[p(\lambda)]$ is equal to
$\prod_{1\leq j<\ell\leq n}(\lambda_{j}-\lambda_{\ell})^{2}.$ (2.3)
Proof
It follows from Definition 2.1 that the discriminant is a polynomial of the coefficients $a_{0},a_{1},\cdots,a_{n-1}$. The coefficients $a_{0},a_{1},\cdots,a_{n-1}$ are elementary symmetric polynomials of the roots $\lambda_{1},\lambda_{2},\cdots,\lambda_{n}$. Therefore the discriminant $\Delta[p(\lambda)]$ is a symmetric polynomial of the roots $\lambda_{1},\lambda_{2},\cdots,\lambda_{n}$. Let us set $\Delta[p(\lambda)]=(-1)^{n(n-1)/2}d(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$.
Now we consider an equality
$\left(\begin{array}[]{cccccccc}1&a_{n-1}&\cdots&\cdots&a_{0}&&&\\ &1&a_{n-1}&\cdots&\cdots&a_{0}&&\\ &&\ddots&\ddots&\ddots&\ddots&\ddots&\\ &&&1&a_{n-1}&\cdots&\cdots&a_{0}\\ n&b_{n-1}&\cdots&b_{2}&b_{1}&&&\\ &n&b_{n-1}&\cdots&b_{2}&b_{1}&&\\ &&\ddots&\ddots&\ddots&\ddots&\ddots&\\ &&&n&b_{n-1}&\cdots&b_{2}&b_{1}\end{array}\right)\left(\begin{array}[]{cccccc}1&&&\lambda_{1}^{2n-2}&\cdots&\lambda_{n}^{2n-2}\\ &\ddots&&\vdots&\ddots&\vdots\\ &&1&\lambda_{1}^{n}&\cdots&\lambda_{n}^{n}\\ &&&\lambda_{1}^{n-1}&\cdots&\lambda_{n}^{n-1}\\ &&&\vdots&\ddots&\vdots\\ &&&\lambda_{1}&\cdots&\lambda_{n}\\ &&&1&\cdots&1\\ \end{array}\right)$ (2.19)
$=$ $\left(\begin{array}[]{cccccccc}1&*&\cdots&*&&&&\\ &1&\ddots&\vdots&&&&\\ &&\ddots&*&&&&\\ &&&1&&&&\\ {}*&*&\cdots&*&\lambda_{1}^{n-1}p^{\prime}(\lambda_{1})&\lambda_{2}^{n-1}p^{\prime}(\lambda_{2})&\cdots&\lambda_{n}^{n-1}p^{\prime}(\lambda_{n})\\ {}*&*&\cdots&*&\lambda_{1}^{n-2}p^{\prime}(\lambda_{1})&\lambda_{2}^{n-2}p^{\prime}(\lambda_{2})&\cdots&\lambda_{n}^{n-2}p^{\prime}(\lambda_{n})\\ {}*&*&\ddots&*&\vdots&\vdots&\ddots&\vdots\\ {}*&*&\cdots&*&p^{\prime}(\lambda_{1})&p^{\prime}(\lambda_{2})&\cdots&p^{\prime}(\lambda_{n})\\ \end{array}\right)$ (2.28)
and compute the determinant of the both sides, we find
$d(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})\prod_{1\leq j<\ell\leq n}(\lambda_{j}-\lambda_{\ell})=\prod_{j=1}^{n}p^{\prime}(\lambda_{j})\prod_{1\leq j<\ell\leq n}(\lambda_{j}-\lambda_{\ell}).$ (2.29)
Dividing the both sides with $\prod_{1\leq j<\ell\leq n}(\lambda_{j}-\lambda_{\ell})$ yields
$d(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ $=$ $\prod_{j=1}^{n}p^{\prime}(\lambda_{j})=\prod_{j=1}^{n}\prod_{\ell(\neq j)}(\lambda_{j}-\lambda_{\ell})$ (2.30)
$=$ $(-1)^{n(n-1)/2}\prod_{1\leq j<\ell\leq n}(\lambda_{j}-\lambda_{\ell})^{2},$
which gives the required result. ∎
Because of Lemma 2.2, a polynomial $p(\lambda)$ has a multiple root, if and only if the discriminant $\Delta[p(\lambda)]$ is zero.
Definition 2.3
Suppose that $K=\mathbb{R}$ or $K=\mathbb{C}$. Let us call a function $f$ from an open set in $K^{d}$ to $\mathbb{C}$ analytic, if there is a Taylor expansion in the form
$f(x)=\sum_{k=0}^{\infty}\sum_{|\alpha|=k}c_{\alpha}(x-a)^{\alpha}$ (2.31)
with $x=(x_{1},x_{2},\cdots,x_{d})$ in the neighborhood of each $a\in K^{d}$ within the domain. Here $\alpha=(\alpha_{1},\alpha_{2}.\cdots,\alpha_{d})$ moves within $\mathbb{Z}_{+}^{d}:=\{\alpha\in\mathbb{Z}^{d}:\alpha_{j}\geq 0,\ j=1,2,\cdots,d\}$, $c_{\alpha}\in\mathbb{C}$ and
$|\alpha|=\alpha_{1}+\alpha_{2}+\cdots\alpha_{d},\ \ \ x^{\alpha}=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{d}^{\alpha_{d}}.$ (2.32)
Moreover we call a function
$F(x)=(f_{1}(x),f_{2},(x),\cdots,f_{n}(x))$ (2.33)
from an open set in $K^{d}$ to $\mathbb{C}^{n}$ analytic, if $f_{1},f_{2},\cdots,f_{n}$ are all analytic.
Supposing that $x_{1},x_{2}.\cdots,x_{d-1}$ are constants, we obtain an analytic function $f_{1}(x)=f(x_{1},x_{2},\cdots,x_{d-1},x)$ of one variable $x\in K$. As is well-known, when $K=\mathbb{C}$, the zero set of an analytic function of one variable is an at most countable set or the whole domain. In the case $K=\mathbb{R}$, the zero set is also an at most countable set or the whole domain, because each of the real and imaginary parts of an analytic function with $K=\mathbb{R}$ can be analytically cotinuated to the complex plane.
[…]
Definition 2.4
Suppose again that $K=\mathbb{R}$ or $K=\mathbb{C}$. If a $K$-valued random variable $X$ satisfies
$\forall x\in K,\ \ \ \mathbb{P}(X=x)=0,$ (2.34)
then $X$ is called a broad continuous random variable (or simply a continuous random variable).
In particular, a random variable with an absolutely continuous measure is broad continuous. A broad continuous random variable is not necessarily absolutely continuous.
In the following, the set of all $N\times N$ matrix on $\mathbb{C}$ is denoted by $\mathbb{M}_{N}(\mathbb{C})$.
Theorem 2.5
Let us put $K=\mathbb{R}$ or $K=\mathbb{C}$. Suppose that $D_{1},D_{2},\cdots,D_{d}$ are open sets on $K$ and $D=D_{1}\times D_{2}\times\cdots\times D_{d}$. We introduce independent and (broad) continuous random variables $X_{1},X_{2},\cdots,X_{d}$: $\Omega\rightarrow K$, and assume that $\mathbb{P}(X_{j}\in D_{j})=1$ for each $j=1,2,\cdots,d$. Let us consider a random matrix $\varphi(X_{1},X_{2},\cdots,X_{d})$ given by an analytic function $\varphi(x_{1},x_{2},\cdots,x_{d})$ ($\varphi:D\rightarrow\mathbb{M}_{N}(\mathbb{C})$). If at least one vector $(x_{1},x_{2},\cdots,x_{d})\in D$ exists and $\varphi(x_{1},x_{2},\cdots,x_{d})$ has $N$ distinct eigenvalues, then the random matrix $\varphi(X_{1},X_{2},\cdots,X_{d})$ has $N$ distinct eigenvalues with probability one.
Remark
If there exits at least one pair of multiple eigenvalues of $\varphi(x_{1},x_{2},\cdots,x_{d})$ for arbitrary $(x_{1},x_{2},\cdots,x_{d})\in D$, then the probability for $\varphi(X_{1},X_{2},\cdots,X_{d})$ to have $N$ distinct eigenvalues is obviously zero. That is, the probability to have $N$ distinct eigenvalues is in general either zero or one.
Proof of Theorem 2.5
We denote the distribution measure of $X_{j}$ by $\mu_{j}$ for $j=1,2,\cdots,d$. Since $X_{j}$ is (broad) continuous, for an at most countable set $A\subset D_{j}$, $\mu_{j}(A)=0$. Let us write the discriminant of the characteristic polynomial of a matrix $P$ as $\Delta(P)$. Because $\Delta(P)$ is a polynomial whose indeterminates are the entries of the matrix $P$, the function $\Delta(\varphi(x_{1},x_{2},\cdots,x_{d}))$ is analytic. As $\Delta(\varphi(x_{1},x_{2},\cdots,x_{d}))$ is analytic, when $x_{1}\in D_{1}$, $x_{2}\in D_{2}$$\cdots,$ $x_{d-1}\in D_{d-1}$ are arbitrarily fixed, the zero set
$\{x\in D_{d}|\Delta(\varphi(x_{1},x_{2},\cdots,x_{d-1},x))=0\}$ (2.35)
is an utmost countable set or the whole $D_{d}$. Therefore we find
$\int_{D_{d}}\ \mathbbm{1}_{\Delta(\varphi(x_{1},\cdots,x_{d-1},x))\neq 0}\ d\mu_{d}(x)$ (2.38)
$=$ $\left\{\begin{array}[]{ll}1,&\exists x_{d},\ \Delta(\varphi(x_{1},\cdots,x_{d-1},x_{d}))\neq 0,\\ 0,&{\rm otherwise}.\end{array}\right$.
Repeatedly using this relation leads to
$\mathbb{P}\left(\varphi(X_{1},X_{2},\cdots,X_{d})\ {\rm has}\ N\ {\rm distinct}\ {\rm eigenvalues}\right)$
$=\int_{D_{1}}d\mu_{1}\int_{D_{2}}d\mu_{2}\cdots\int_{D_{d}}d\mu_{d}\ \mathbbm{1}_{\Delta(\varphi)\neq 0}$
$=\left\{\begin{array}[]{ll}1,&\exists\{x_{1},x_{2},\cdots,x_{d}\},\ \Delta(\varphi(x_{1},x_{2},\cdots,x_{d}))\neq 0,\\ 0,&{\rm otherwise},\end{array}\right$. (2.41)
which gives the required result. ∎
Example 2.6
Let us put $\mathbbm{i}=\sqrt{-1}$. The whole set of $N\times N$ hermitian matrices is identified with $\mathbb{R}^{N^{2}}$ by means of the following map $\varphi:\mathbb{R}^{N^{2}}\rightarrow\mathbb{M}_{N}(\mathbb{C})$.
$\phi(x_{11},\cdots,x_{NN})$ (2.50)
$=$ $\left(\begin{array}[]{cccc}x_{11}&x_{12}&\cdots&x_{1N}\\ x_{12}&x_{22}&\cdots&x_{2N}\\ \vdots&\vdots&\ddots&\vdots\\ x_{1N}&x_{2N}&\cdots&x_{NN}\end{array}\right)+\mathbbm{i}\left(\begin{array}[]{ccccc}0&-x_{21}&\cdots&&-x_{N1}\\ x_{21}&0&\cdots&&-x_{N2}\\ \vdots&\vdots&\ddots&&\vdots\\ x_{N1}&x_{N2}&\cdots&&0\end{array}\right).$
We define the subset $S\subset\mathbb{R}^{N^{2}}$ as the whole of the $N\times N$ hermitian matrices with multiple eigenvalues. Then one can prove that the Lebesgue measure $\mathfrak{m}(S)$ is zero[5]. This fact can also be proved by using Theorem 2.5, as explained below.
Let us arbitrarily fix $n\in\mathbb{N}$ and set $I_{n}=[-n,n]$. If each of independent random variables $X_{11},\cdots,X_{NN}$ is distributed according to a probability density function
$p(x)=\left\{\begin{array}[]{ll}1/(2n),&x\in I_{n},\\ 0,&x\notin I_{n},\end{array}\right$. (2.52)
the probability for $\varphi(X_{11},\cdots,X_{NN})$ to have multiple eigenvalues is equal to $(2n)^{-N^{2}}\mathfrak{m}(S\cap I_{n}^{N^{2}})$. On the other hand, if we put $x_{j\ell}=j\delta_{j\ell}$ ($\delta_{j\ell}$ is the Kronecker delta), the eigenvalues of $\varphi(x_{11},\cdots,x_{NN})$ do not degenerate. Then it follows from Theorem 2.5 that the probability for the eigenvalues of $\varphi(X_{11},\cdots,X_{NN})$ to degenerate is zero. Therefore $\mathfrak{m}(S\cap I_{n}^{N^{2}})=0$ holds for arbitrary $n\in\mathbb{N}$ and we obtain
$\mathfrak{m}(S)=\mathfrak{m}\left(\bigcup_{n=1}^{\infty}\left(S\cap I_{n}^{N^{2}}\right)\right)=\lim_{n\rightarrow\infty}\mathfrak{m}\left(S\cap I_{n}^{N^{2}}\right)=0.$ (2.53)
Example 2.7
Suppose that $E_{j\ell}\in\mathbb{M}_{N}(\mathbb{C})$ is a matrix unit, so that the $(j,\ell)$ entry of $E_{j\ell}$ is one and the other entries are all zero. We define
$P_{j}=E_{jj},\ \ \ Q_{j\ell}=-\mathbbm{i}E_{j\ell}+\mathbbm{i}E_{\ell j}$ (2.54)
and an analytic function $\varphi:\mathbb{R}^{N^{2}}\rightarrow\mathbb{M}_{N}(\mathbb{C})$ as
$\varphi(x_{11},\cdots,x_{NN})$ (2.55)
$=$ $\left(\prod_{j=1}^{N-1}\prod_{\ell=j+1}^{N}{\rm exp}(\mathbbm{i}x_{\ell j}P_{\ell})\ {\rm exp}(\mathbbm{i}x_{j\ell}Q_{j\ell})\right)\ \left(\prod_{j=1}^{N}{\rm exp}(\mathbbm{i}x_{jj}P_{j})\right),$
where the order of a product of matrices is defined as $\prod_{j=1}^{n}A_{j}=A_{1}A_{2}\cdots A_{n}$. If the probability density function $p_{11},\cdots,p_{NN}$ for each of independent $\mathbb{R}$-valued random variables $X_{11},\cdots,X_{NN}$ is
$p_{j\ell}(x)=$ $1/(2\pi),$ $x\in[0,2\pi]\ {\rm for}\ j\geq\ell,$
$p_{j\ell}(x)=$ $2(\ell-j)\ \sin x\ \cos^{2(\ell-j)-1}x,$ $x\in[0,\pi/2]\ {\rm for}\ j<\ell,$ (2.56)
then the probability measure of $\varphi(X_{11},\cdots,X_{NN})$ is the Haar measure of the unitary group[6]. In a special case $x_{j\ell}=2\pi j\delta_{j\ell}/N$, the eigenvalues of $\varphi(x_{11},\cdots,x_{NN})$ do not degenerate. Therefore we can see from Theorem 2.5 that the probability for the eigenvalues of $\varphi(X_{11},\cdots,X_{NN})$ to degenerate is zero.
Lemma 2.8
We set $K=\mathbb{R}$ or $K=\mathbb{C}$. Let us consider a $K$-linear map $\varphi:K^{d}\rightarrow\mathbb{M}_{N}(\mathbb{C})$. Suppose that a set of constants $c_{1},\cdots,c_{r}\in K$ exists for $r<d$, so that $\varphi(c_{1},\cdots,c_{r},x_{r+1},\cdots,x_{d})$ has at least one pair of multiple eigenvalues for arbitrary $x_{r+1},\cdots,x_{d}$. Then $\varphi(0,\cdots,0,x_{r+1},\cdots,x_{d})$ (the constant arguments $c_{1},\cdots,c_{r}$ of $\varphi$ are replaced with $0,\cdots,0$) has at least one pair of multiple eigenvalues.
Proof
We introduce notations
$y=(y_{1},\cdots,y_{r}),\ \ \ {\tilde{x}}=(x_{r+1},\cdots,x_{d})$ (2.57)
and denote the discriminant of the characteristic polynomial of a matrix $P$ by $\Delta(P)$. Because $\varphi(x_{1},\cdots,x_{d})$ is a linear homogeneous polynomial of $x_{1},\cdots,x_{d}$ and $\Delta(P)$ is a degree N(N1)N(N-1) homogeneous polynomial of the entries of $P$, we can write $)\Delta(\varphi(y,{\tilde{x}}$)) in the form
$\Delta(\varphi(y,{\tilde{x}}))=\sum_{|\alpha|\leq N(N-1)}\sum_{|\beta|=N(N-1)-|\alpha|}\kappa_{\beta\alpha}y^{\beta}{\tilde{x}}^{\alpha},\ \ \ \kappa_{\beta\alpha}\in\mathbb{C}.$ (2.58)
The multiple index notation was introduced in Definition 2.3. When $y=c=(c_{1},\cdots,c_{r})$, we assume that $\Delta(\varphi(c,{\tilde{x}}))=0$ for arbitrary $\tilde{x}$}. Therefore
$\sum_{|\beta|=N(N-1)-|\alpha|}\kappa_{\beta\alpha}c^{\beta}=0$ (2.59)
for arbitrary $\alpha$ with $|\alpha|\leq N(N-1)$. In a special case $|\alpha|=N(N-1)$, we obtain $\kappa_{(0,\cdots,0)\alpha}=0$. Then it follows that
$)\displaystyle\Delta(\varphi(0,\cdots,0,{\tilde{x}}$)) $=$ $\sum_{|\alpha|=N(N-1)}\kappa_{(0,\cdots,0)\alpha}{\tilde{x}}^{\alpha}$ (2.60)
$+\sum_{|\alpha|<N(N-1)}\sum_{|\beta|=N(N-1)-|\alpha|}\kappa_{\beta\alpha}(0,\cdots,0)^{\beta}{\tilde{x}}^{\alpha}$
$=$ $0,$
which gives the required result. ∎
Using Lemma 2.8, we are able to prove the following Corollary 2.9, which generalizes Theorem 2.5 for a linear map $\varphi$ on $\mathbb{R}^{d}$. It gives a sufficient condition for a random matrix $\varphi(X_{1},\cdots,X_{d})$ to have $N$ distinct eigenvalues, when some of $X_{1},\cdots,X_{d}$ are not necessarily continuous.
Corollary 2.9
Let $K=\mathbb{R}$ or $K=\mathbb{C}$. Suppose that $\varphi:K^{d}\rightarrow\mathbb{M}_{N}(\mathbb{C})$ is a $K$-linear map. Moreover we assume that a vector $(x_{r+1},\cdots,x_{d})\in K^{d-r}$ exists for $r<d$, so that $\varphi(0,\cdots,0,x_{r+1},\cdots,x_{d})$ has $N$ distinct eigenvalues. If $X_{1},\cdots,X_{d}:\Omega\rightarrow K$ are independent random variables and $X_{r+1},\cdots,X_{d}$ are (broad) continuous, then $\varphi(X_{1},\cdots,X_{d})$ have $N$ distinct eigenvalues with probability one.
Proof
The contraposition of Lemma 2.8 implies that $\varphi(c_{1},\cdots,c_{r},x_{r+1},\cdots,x_{d})$ has $N$ distinct eigenvalues for arbitrary $c_{1},\cdots,c_{r}\in K$. Then Theorem 2.5 gives the required result, because $X_{r+1},\cdots,X_{d}$ are (broad) continuous.
3 Eigenvalue degeneracy and perfect matching of graphs
3.1 Asymmetric matrices
Let us first consider asymmetric matrices and corresponding bipartite graphs. In particular, we focus on the relation between eigenvalue degeneracy and perfect matching of graphs. In the following, $E_{j\ell}$ stands for a matrix unit defined in Example 2.7.
Lemma 3.1
Let us suppose that $\sigma$ is an arbitrary element of the $N$-th order symmetric group $\mathfrak{S}_{N}$ and define an $N\times N$ matrix
$\varphi(x_{1},x_{2},\cdots,x_{N})=\sum_{j=1}^{N}x_{j}E_{j\ \sigma(j)},\ \ \ x_{j}\in K,$ (3.1)
where $K$ is $\mathbb{R}$ or $\mathbb{C}$. That is, the $(m,n)$ entry $\varphi_{mn}$ of $\varphi(x_{1},x_{2},\cdots,x_{N})$ is
$\varphi_{mn}=\sum_{j=1}^{N}x_{j}\ \delta_{jm}\ \delta_{\sigma(j)n}=x_{m}\delta_{\sigma(m)n}.$ (3.2)
Then a set $\{x_{1},x_{2},\cdots,x_{N}\}$ exists, for which the $N$ eigenvalues of $\varphi$ are different from each other and all non-zero.
Proof
A permutation $\sigma\in\mathfrak{S}_{N}$ can be written as a product of cyclic permutations involving no common indices. That is, denoting a cyclic permutation
$n_{1}\rightarrow n_{2}\rightarrow\cdots\rightarrow n_{\ell}\rightarrow n_{1}$
as $(n_{1}n_{2}\cdots n_{\ell})$, we can write $\sigma$ in the form
$\sigma=(n_{11}n_{12}\cdots n_{1\ell_{1}})(n_{21},n_{22},\cdots,n_{2\ell_{2}})\cdots(n_{r1}n_{r2}\cdots n_{r\ell_{r}}),$ (3.3)
where $1,2,\cdots,N$ is reordered to
$n_{11},n_{12},\cdots,n_{1\ell_{1}},n_{21},n_{22},\cdots,n_{2\ell_{2}}.\cdots,n_{r1},n_{r2},\cdots,n_{r\ell_{r}}.$ (3.4)
Now the characteristic polynomial of $\varphi$, which is the determinant of $\lambda I_{N}-\varphi$ with the $N\times N$ identity matrix $I_{N}$, can be factored as
$\prod_{j=1}^{r}\left(\lambda^{\ell_{j}}-x_{n_{j1}}x_{n_{j2}}\cdots x_{n_{j\ell_{j}}}\right).$ (3.5)
Here each factor corresponds to a cyclic permutation $(n_{j1}n_{j2}\cdots n_{j\ell_{j}})$. If we choose $x_{1},x_{2},\cdots,x_{N}$ such that
$\sqrt[\ell_{j}]{\left|x_{n_{j1}}x_{n_{j2}}\cdots x_{n_{j\ell_{j}}}\right|},\ \ \ j=1,2,\cdots r$
are different from each order and all non-zero, we obtain the required result. ∎
Now we introduce a terminology of graph theory.
Definition 3.2
[…]
$(v_{1},v_{2}),(v^{\prime}_{1},v^{\prime}_{2})\in M\ {\rm and}\ (v_{1},v_{2})\neq(v^{\prime}_{1},v^{\prime}_{2})\Longrightarrow v_{1}\neq v^{\prime}_{1}\ {\rm and}\ v_{2}\neq v^{\prime}_{2},$ (3.6)
then $M$ is a matching of $G$. Moreover, if a matching $M$ meets a condition that $v_{2}\in{\cal V}_{2}$ satisfying $(v_{1},v_{2})\in M$ exists for an arbitrary $v_{1}\in{\cal V}_{1}$, then $M$ is a perfect matching of $G$. In that case $G$ is said to have a perfect matching $M$.
In the following, we consider a bipartite graph $G=({\cal V}^{R},{\cal V}^{C},{\cal E})$, supposing that
${\cal V}^{R}=\{v_{1}^{R},v_{2}^{R},\cdots,v_{N}^{R}\},\ \ \ {\cal V}^{C}=\{v_{1}^{C},v_{2}^{C},\cdots,v_{N}^{C}\},\ \ \ {\cal E}\subset{\cal V}^{R}\times{\cal V}^{C}.$ (3.7)
[…]
$\Lambda\subset{\cal I}_{N}\times{\cal I}_{N}\longleftrightarrow{\cal E}=\{(v_{j}^{R},v_{\ell}^{C})|(j,\ell)\in\Lambda\}\subset{\cal V}^{R}\times{\cal V}^{C}.$ (3.8)
Hereafter we abbreviate $(v_{j}^{R},v_{\ell}^{C})\in{\cal E}$ as $\langle j,\ell\rangle$. We moreover write a family of parameters $x_{j\ell}$ with $(j,\ell)\in\Lambda$ as $x_{\Lambda}$.
Lemma 3.3
Let us suppose that $\Lambda\subset{\cal I}_{N}\times{\cal I}_{N}$ and $\varphi(x_{\Lambda})=\sum_{(j,\ell)\in\Lambda}x_{j\ell}E_{j\ell}$ with $x_{j\ell}\in K$ ($K$ is $\mathbb{R}$ or $\mathbb{C}$). If a bipartite graph $G=({\cal V}^{R},{\cal V}^{C},{\cal E})$ has a perfect matching, then an $x_{\Lambda}$ exits, so that all the $N$ eigenvalues of $\varphi(x_{\Lambda})$ are different from each other and all non-zero.
Proof
[…]
${\cal M}=\{\langle 1,\sigma(1)\rangle,\langle 2,\sigma(2)\rangle,\cdots,\langle N,\sigma(N)\rangle\}.$ (3.9)
Correspondingly, we also define
$M=\{(1,\sigma(1)),(2,\sigma(2)),\cdots,(N,\sigma(N))\}.$ (3.10)
Then, supposing that $x_{j\ell}=0$ for $(j,\ell)\in\Lambda\setminus M$, we apply Lemma 3.1 to
${\tilde{\varphi}}(x_{M})=\varphi(x_{\Lambda\setminus M}=0,x_{M})=\sum_{j=1}^{N}x_{j\sigma(j)}E_{j\sigma(j)}$ (3.11)
and obtain the required result. ∎
Corresponding to a bipartite graph $G=({\cal V}^{R},{\cal V}^{C},{\cal E})$ and $S\subset{\cal I}_{N}$, we define a bipartite subgraph $G[S]=\left({\cal V}^{R}[S],{\cal V}^{C}[S],{\cal E}[S]\right)$ as
${\cal V}^{R}[S]=\{v_{j}^{R}|j\in S\},\ \ \ {\cal V}^{C}[S]=\{v_{j}^{C}|j\in S\},\ \ \ {\cal E}[S]=\{\langle j,\ell\rangle\in{\cal E}|j,\ell\in S\}.$ (3.12)
Theorem 3.4
Let us suppose that $\Lambda\subset{\cal I}_{N}\times{\cal I}_{N}$ and $\varphi(x_{\Lambda})=\sum_{(j,\ell)\in\Lambda}x_{j\ell}E_{j\ell}$ with $x_{j\ell}\in K$. Here $K$ is $\mathbb{R}$ or $\mathbb{C}$. Then the following conditions are equivalent.
(1) An $x_{\Lambda}$ exists, so that the $N$ eigenvalues of $\varphi(x_{\Lambda})$ are different from each other.
(2) An $x_{\Lambda}$ exists, so that $\varphi(x_{\Lambda})$ has no zero eigenvalues or a simple zero eigenvalue. Here a simple zero eigenvalue means an eigenvalue with multiplicity one.
(3) The bipartite graph $G=({\cal V}^{R},{\cal V}^{C},{\cal E})$ has a perfect matching, or $j\in{\cal I}_{N}$ exists and the bipartite subgraph $G[{\cal I}_{N}\setminus\{j\}]$ has a perfect matching.
Proof
(1) $\Rightarrow$ (2): It is obvious.
(2) $\Rightarrow$ (3): If $\varphi(x_{\Lambda})$ has no zero eigenvalues, then
$\det\varphi(x_{\Lambda})=\det\left(\sum_{(j,\ell)\in\Lambda}x_{j\ell}E_{j\ell}\right)=\sum_{\sigma\in\mathfrak{S}_{N}}{\rm sign}(\sigma)\prod_{n=1}^{N}\left(\sum_{(j,\ell)\in\Lambda}x_{j\ell}\delta_{jn}\delta_{\ell\sigma(n)}\right)\neq 0.$ (3.13)
Therefore there exists at least one $\sigma\in\mathfrak{S}_{N}$ satisfying $(n,\sigma(n))\in\Lambda$ for all $n=1,2,\cdots,N$. This means that $G$ has a perfect matching.
Let us next consider the characteristic polynomial $\varpi(\lambda)=\det(\lambda I_{N}-\varphi(x_{\Lambda})$. If $\varphi(x_{\Lambda})$ has a simple zero eigenvalue, $\varpi^{\prime}(0)\neq 0$. Then at least one of the order N1N-1 principal minors of $\varphi(x_{\Lambda})$ is non-zero, since the sum of those principal minors is $\varpi^{\prime}(0)$. Then it follows that $j\in{\cal I}_{N}$ exists and $G[{\cal I}_{N}\setminus\{j\}]$ has a perfect matching.
(3) $\Rightarrow$ (1): When $G$ has a perfect matching, it is obvious because of Lemma 3.3. When $G[{\cal I}_{N}\setminus\{j\}]$ has a perfect matching for $j\in{\cal I}_{N}$, let us consider the $(N-1)\times(N-1)$ submatrix generated from $\varphi(x_{\Lambda})$ by removing the $j$-th row and $j$-th column. We can see from Lemma 3.3 that there exists a family of parameters $x_{k\ell}$ with $\langle k,\ell\rangle\in{\cal E}[{\cal I}_{N}\setminus\{j\}]$, so that the N1N-1 submatrix eigenvalues are different from each other and all non-zero. Moreover we put $x_{k\ell}=0$ when $\langle k,\ell\rangle\in{\cal E}\setminus{\cal E}[{\cal I}_{N}\setminus\{j\}]$. Then the $N$ eigenvalues of $\varphi(x_{\Lambda})$ are different from each other, because they consist of a simple zero eigenvalue and the submatrix eigenvalues. ∎
As an application, let us prove the following theorem.
Theorem 3.5
Suppose that $N\geq 2$, $\Lambda\subset{\cal I}_{N}\times{\cal I}_{N}$ and $\varphi(x_{\Lambda})=\sum_{(j,\ell)\in\Lambda}x_{j\ell}E_{j\ell}$ with $x_{jl}\in K$. Here $K$ is $\mathbb{R}$ or $\mathbb{C}$. If $\Lambda\geq N^{2}-2N+2$, there exits an $x_{\Lambda}$, so that the $N$ eigenvalues of $\varphi(x_{\Lambda})$ are different from each other. Here the number of the elements of a set $A$ is denoted by #A\#A.
If $\Lambda<N^{2}-2N+2$, there are counterexamples. For example, if we have
$\Lambda=\{(j,\ell)|j\geq 3\}\cup\{(2,1)\},$ (3.14)
then $\Lambda=N^{2}-2N+1$ and the zero eigenvalues of $\varphi(x_{\Lambda})$ degenerate.
In order to prove Theorem 3.5, we use Hall’s marriage theorem in graph theory. For a bipartite graph $G=({\cal V}_{1},{\cal V}_{2},{\cal E})$, a subset $J_{1}\subset{\cal V}_{1}$ and a subset $J_{2}\subset{\cal V}_{2}$, we define
$\Gamma_{G}(J_{1})$ $=$ $\{v_{2}\in{\cal V}_{2}|\exists v_{1}\in J_{1},(v_{1},v_{2})\in{\cal E}\},$
$\Gamma_{G}(J_{2})$ $=$ $\{v_{1}\in{\cal V}_{1}|\exists v_{2}\in J_{2},(v_{1},v_{2})\in{\cal E}\}.$ (3.15)
If $J_{1}$ or $J_{2}$ is a one-point set $\{v\}$, then $\Gamma_{G}(\{v\})$ is simply written as $\Gamma_{G}(v)$.
Theorem 3.6 (Hall’s marriage theorem)
Let ${\cal V}_{1}$ be a finite set. For a bipartite graph $G=({\cal V}_{1},{\cal V}_{2},{\cal E})$, the following conditions are equivalent.
(a) A perfect matching exists.
(b) For an arbitrary $A\subset{\cal V}_{1}$, $A\leq\#\Gamma_{G}(A)$.
This (b) is called Hall’s condition. This theorem is named after Hall who first published it in 1935[8] and a proof of it can be found in [7] (Theorem 1.1.3). It leads to the following lemma.
Lemma 3.7
Suppose that $G=({\cal V}_{1},{\cal V}_{2},{\cal E})$ is a bipartite graph with ${\cal V}_{1}=\#{\cal V}_{2}=n$. If ${\cal E}\geq n^{2}-n+1$, $G$ has a perfect matching.
Proof
Let us arbitrarily take $A\subset{\cal V}_{1}$ ($A\neq\emptyset$). If $v\in A$ and $w\in{\cal V}_{2}$ satisfies $w\notin\Gamma_{G}(A)$, then $(v,w)\notin{\cal E}$. Hence
$A\times\left({\cal V}_{2}\setminus\Gamma_{G}(A)\right)\subset\left({\cal V}_{1}\times{\cal V}_{2}\right)\setminus{\cal E}.$ (3.16)
Therefore
$\left(\#A\right)\left(n-\#\Gamma_{G}(A)\right)\leq n^{2}-\#{\cal E}\leq n-1.$ (3.17)
If follows that
$\#\Gamma_{G}(A)$ $\geq$ $n-\frac{n-1}{\#A}=(\#A-1)\frac{n}{\#A}+\frac{1}{\#A}\geq\#A-1+\frac{1}{\#A}$ (3.18)
$>$ $\#A-1.$
Consequently Hall’s condition $\Gamma_{G}(A)\geq\#A$ holds. ∎
Based on the above preparations, we prove Theorem 3.5.
Proof of Theorem 3.5
Let us show that the condition (3) of Theorem 3.4 holds. We first suppose that the bipartite graph $G=({\cal V}^{R},{\cal V}^{C},{\cal E})$ has no perfect matchings. According to Hall’s marriage theorem, $A\subset{\cal V}^{R}$ exits, so that $A>\#\Gamma_{G}(A)$. Then
${\cal K}=\{j|v^{R}_{j}\in A\ {\rm and}\ v^{C}_{j}\notin\Gamma_{G}(A)\}$ (3.19)
is not empty and we can take $k\in{\cal K}$. The definition of $\cal K$} means
$\langle k,\ell\rangle\in{\cal E}\Longrightarrow v_{\ell}\in\Gamma_{G}(A)$ (3.20)
and
$\langle j,k\rangle\in{\cal E}\Longrightarrow v_{j}\in{\cal V}^{R}\setminus A.$ (3.21)
On the other hand, it follows from
${\cal E}[{\cal I}_{N}\setminus\{k\}]=\{\langle j,\ell\rangle|j\neq k\ {\rm and}\ \ell\neq k\}$ (3.22)
that
${\cal E}\setminus{\cal E}[{\cal I}_{N}\setminus\{k\}]=\{\langle j,\ell\rangle|j=k\ {\rm or}\ \ell=k\}.$ (3.23)
Putting (3.20), (3.21) and (3.23) together, we obtain
${\cal E}\setminus{\cal E}[{\cal I}_{N}\setminus\{k\}]\subset\{\langle j,k\rangle|v_{j}\in{\cal V}^{R}\setminus A\}\cup\{\langle k,\ell\rangle|v_{\ell}\in\Gamma_{G}(A)\},$ (3.24)
which yields
${\cal E}-\#{\cal E}[{\cal I}_{N}\setminus\{k\}]\leq N-\#A+\#\Gamma_{G}(A)\leq N-1.$ (3.25)
Let us now suppose that ${\cal E}=\#\Lambda\geq N^{2}-2N+2$. Then one can see that
${\cal E}[{\cal I}_{N}\setminus\{k\}]\geq\#{\cal E}-N+1\geq N^{2}-3N+3=(N-1)^{2}-(N-1)+1.$ (3.26)
On account of Lemma 3.7, $G[{\cal I}_{N}\setminus\{k\}]$ has a perfect marching. Therefore the condition (3) of Theorem 3.4 holds. ∎
The following corollary of Theorem 3.5 gives the maximum number of continuous random matrix entries, when we have eigenvalue degeneracy.
Corollary 3.8
We set $K=\mathbb{R}$ or $K=\mathbb{C}$. Suppose that $N\geq 2$ and let $X_{j\ell}:\Omega\rightarrow K$ ($j,\ell=1,2,\cdots,N$) be independent random variables. We consider an $N\times N$ square random matrix $(X_{j\ell})_{j,\ell=1,2,\cdots,N}$. If more than or equal to $N^{2}-2N+2$ of the random variables $X_{j\ell}$ are (broad) continuous, the random matrix $(X_{j\ell})$ has $N$ different eigenvalues with probability one.
Proof
It immediately follows from Corollary 2.9 and Theorem 3.5.
3.2 Symmetric matrices
Next we consider symmetric matrices and corresponding graphs. In order to treat symmetric matrices, we introduce
${\cal I}^{(2)}_{N}=\{\{j,\ell\}|j,\ell\in{\cal I}_{N}\}.$ (3.27)
Moreover, for ${\bf e}=\{j,\ell\}\in{\cal I}^{(2)}_{N}$, we define
$E_{\bf e}=E_{\{j,\ell\}}=\frac{1}{2}(E_{j\ell}+E_{\ell j}).$ (3.28)
Lemma 3.9
Let us take an element $\sigma$ of the $N$-th order symmetric group $\mathfrak{S}_{N}$ and introduce the notations
${\bf S}=\left\{\{j,\sigma(j)\}|j\in{\cal I}_{N}\right\}$ (3.29)
and
$\varphi(x_{\bf S})=\sum_{{\bf e}\in{\bf S}}x_{\bf e}E_{\bf e}.$ (3.30)
Suppose that $K$ is $\mathbb{R}$ or $\mathbb{C}$. Then, for an $x_{S}\in K^{\#{\bf S}}$, the eigenvalues of $\varphi$ do not degenerate, and are all different from zero.
Proof
In the same way as the proof of Lemma 3.1, by decomposing $\sigma$ into a product of cyclic permutations, we find that it is sufficient to show this lemma when $\sigma$ is a cyclic permutation $(12\cdots N)$. Let us restrict ourselves to that case.
If $N\leq 2$, $\bf S$} consists of a unique element $\{1,\sigma(1)\}$. Then the required result holds when $x_{\{1,\sigma(1)\}}\neq 0$. In the following, we assume $N\geq 3$. Then the number of the elements of $\bf S$} is $N$. We write the variable $x_{\{j,\sigma(j)\}}$ as $x_{j}$. Let us put $x_{1}=x_{2}=\cdots=x_{N-1}=1$ and $x_{N}=\epsilon$, and consider the matrix $P(\epsilon)=\varphi(1,\cdots,1,\epsilon)$. When $\epsilon=0$, $P(0)$ is a tridiagonal matrix and the eigenvalues are $\cos\left(k\pi/(N+1)\right)$ ($k=1,2,\cdots,N)$. The required result evidently holds for an even $N$. In the following we consider the odd $N$ case. As the eigenvalues of $P(0)$ do not degenerate, the eigenvalues of $P(\epsilon)$ do not degenerate either, provided that $|\epsilon|$ is small enough, owing to the continuity of eigenvalues. On the other hand, since $\det P(\epsilon)=\epsilon/2^{N}$, all the eigenvalues are non-zero, if $\epsilon\neq 0$. Therefore the require result holds, when $|\epsilon|$ has a non-zero but sufficiently small value. ∎
By using Lemma 3.9, similarly as Lemma 3.3 and Theorem 3.4, we can prove the following lemma and theorem for symmetric matrices.
Lemma 3.10
Let us suppose that $\mathfrak{L}\subset{\cal I}^{(2)}_{N}$ and $\varphi(x_{\mathfrak{L}})=\sum_{{\bf e}\in\mathfrak{L}}x_{\bf e}E_{\bf e}$ with $x_{\bf e}\in K$ ($K$ is $\mathbb{R}$ or $\mathbb{C}$). Moreover we define
$\mathfrak{E}=\Big\{\langle j,\ell\rangle\in{\cal V}^{R}\times{\cal V}^{C}\Big|\{j,\ell\}\in\mathfrak{L}\Big\}.$ (3.31)
If a bipartite graph $G=({\cal V}^{R},{\cal V}^{C},\mathfrak{E})$ has a perfect matching, then an $x_{\mathfrak{L}}$ exits, so that all the $N$ eigenvalues of $\varphi(x_{\mathfrak{L}})$ are different from each other and all non-zero.
Let $S$ be a subset of ${\cal I}_{N}$. Corresponding to a bipartite graph $G=({\cal V}^{R},{\cal V}^{C},\mathfrak{E})$, a bipartite subgraph $G[S]=\left({\cal V}^{R}[S],{\cal V}^{C}[S],\mathfrak{E}[S]\right)$ is defined as
${\cal V}^{R}[S]=\{v_{j}^{R}|j\in S\},\ \ \ {\cal V}^{C}[S]=\{v_{j}^{C}|j\in S\},\ \ \ \mathfrak{E}[S]=\{\langle j,\ell\rangle\in\mathfrak{E}|j,\ell\in S\}.$ (3.32)
Theorem 3.11
Let us suppose that $\mathfrak{L}\subset{\cal I}^{(2)}_{N}$ and $\varphi(x_{\mathfrak{L}})=\sum_{{\bf e}\in{\cal W}}x_{\bf e}E_{\bf e}$ with $x_{\bf e}\in K$ ($K$ is $\mathbb{R}$ or $\mathbb{C}$). Then the following conditions are equivalent.
(1) An $x_{\mathfrak{L}}$ exists, so that the $N$ eigenvalues of $\varphi(x_{\mathfrak{L}})$ are different from each other.
(2) An $x_{\mathfrak{L}}$ exists, so that $\varphi(x_{\mathfrak{L}})$ has no zero eigenvalues or a simple zero eigenvalue.
(3) The bipartite graph $G=({\cal V}^{R},{\cal V}^{C},\mathfrak{E})$ has a perfect matching, or $j\in{\cal I}_{N}$ exists and the bipartite subgraph $G[{\cal I}_{N}\setminus\{j\}]$ has a perfect matching.
4 Eigenvalue degeneracy probability for random matrices
Let $K=\mathbb{R}$ or $K=\mathbb{C}$. We suppose that $Z_{j\ell}$ ($j,\ell=1,2,\cdots,N)$ are independent and (broad) continuous $K$-valued random variables. Moreover we introduce independent random variables $Y_{j\ell}$ ($j,\ell=1,2,\cdots,N$) obeying Bernoulli distribution with a parameter $p$. A random variable obeying Bernoulli distribution with a parameter $p$ takes a value $1$ with probability $p$ and another value 0 with probability 1p1-p.
We consider the probability for the random matrix $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$ to have $N$ distinct eigenvalues. First let us arbitrarily fix $y_{j\ell}\in\{0,1\}$ ($j,\ell=1,2,\cdots,N$) and evaluate the conditional probability under the condition $Y_{j\ell}=y_{j\ell}$. It follows from $\bf Theorem2.5$} that this conditional probability is zero or one, and the necessary and sufficient condition to have the conditional probability one is:
there exits at least one matrix $(z_{j\ell})_{j,\ell=1,2,\cdots,N}$ with $z_{j\ell}\in K$, so that the matrix $(y_{j\ell}z_{j\ell})_{j,\ell=1,2,\cdots,N}$ has $N$ distinct eigenvalues.
This conditional probability does not depend on the distributions of $Z_{j\ell}$, provided that they are independent and (broad) continuous. Therefore the probability for the matrix $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$ to have $N$ distinct eigenvalues does not depend on the distributions of $Z_{j\ell}$ either. It turns out to depend only on $N$ and $p$. Moreover, we can readily see from Theorem 3.4 that this probability is equal to the probability for the random bipartite graph $G=({\cal V}^{R},{\cal V}^{C},{\cal E})$, ${\cal E}=\{\langle j,\ell\rangle|Y_{j\ell}=1\}$ to satisfy the following condition.
Condition 4.1
The bipartite graph $G=({\cal V}^{R},{\cal V}^{C},{\cal E})$ has a perfect matching, or $j\in{\cal I}_{N}$ exists and the bipartite subgraph $G[{\cal I}_{N}\setminus\{j\}]$ has a perfect matching. Here ${\cal I}_{N}=\{1,2,\cdots,N\}$.
Thus, as a simpler problem, let us first consider the probability for a random graph to have a perfect matching. For that problem, there is a known result explained below as Theorem 4.2.
Remark
Another consequence of Theorem 3.4 is that the probability for the matrix $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$ to have $N$ distinct eigenvalues is equal to the probability that it has no zero eigenvalues or a simple zero eigenvalue. On the other hand, if an $N\times N$ matrix has $N$ distinct eigenvalues, then in general it has no zero eigenvalues or a simple zero eigenvalue. Therefore, if the eigenvalues of $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$ degenerate, then zero eigenvalues exist and degenerate with probability one. Namely, as for this random matrix model, it can be said that accumulation of the eigenvalues to the origin causes eigenvalue degeneracy.
4.1 The probability for a random graph to have a perfect matching
In this subsection, we introduce the following theorem for the asymptotic probability for a random bipartite graph to have a perfect matching. This theorem was originally proved by Erdös and Rényi[9]. It is useful in the study of asymptotic degeneracy probability for the eigenvalues of the corresponding random matrices,
Theorem 4.2
We consider a random bipartite graph $G_{n}=({\cal V}_{1},{\cal V}_{2},{\cal E})$ with an edge set ${\cal E}\subset{\cal V}_{1}\times{\cal V}_{2}$ and ${\cal V}_{1}=\#{\cal V}_{2}=n$. Here vertex sets are
${\cal V}_{1}=\{v_{1}v_{2},\cdots,v_{n}\},\ \ \ {\cal V}_{2}=\{w_{1},w_{2},\cdots,w_{n}\}$ (4.1)
and $(v_{j},w_{\ell})\in{\cal V}_{1}\times{\cal V}_{2}$ is abbreviated as $\langle j,\ell\rangle$. Each $\langle j,\ell\rangle$ ($j,\ell=1,2,\cdots,n$) independently satisfies $\langle j,\ell\rangle\in{\cal E}$ with probability $p$. Suppose that the probability parameter $p$ is dependent on $n$ and written as
$p=p(n)=\frac{\log n+c}{n}+o\left(\frac{1}{n}\right),\ \ \ n\rightarrow\infty$ (4.2)
with a constant $c\in\mathbb{R}$. Then the probability $\mathbb{P}$ for the bipartite graph $G_{n}$ to have a perfect matching is asymptotically evaluated as
$\lim_{n\rightarrow\infty}\mathbb{P}(G_{n}\ {\rm has}\ {\rm a}\ {\rm perfect}\ {\rm matching})=e^{-\lambda},$ (4.3)
where $\lambda=2e^{-c}$.
In the remaining part of this subsection, we give a step-by-step proof of Theorem 4.2 in reference to Frieze and Karoński’s book[10]. Intermediate results are also useful in the study of eigenvalue degeneracy. Let us begin with a definition and a lemma.
Definition 4.3
Suppose that $\mathbb{Z}_{+}=\{m\in\mathbb{Z}|m\geq 0\}$ and $k\in\mathbb{Z}_{+}$. We define a $k$-th order polynomial $\binom{x}{k}$ with indeterminates $x$ as
$\binom{x}{0}=1\ \ {\rm and}\ \ \binom{x}{k}=\frac{x(x-1)\cdots(x-k+1)}{k!}\ {\rm for}\ k\geq 1.$ (4.4)
When $x=n\in\mathbb{Z}+$, $\dbinom{n}{k}\geq 0$. In particular, if $n<k$, we have $\dbinom{n}{k}=0$.
Lemma 4.4
For $k,m,n\in\mathbb{Z}_{+}$, the following equality holds.
$\sum_{k=0}^{m}(-1)^{k}\binom{n+1}{k}=(-1)^{m}\binom{n}{m}$. (4.5)
Proof
Mathematical induction with respect to $m$ is employed. For $m=0$, both sides of the equality is $1$. Let us assume that the equality holds for general $m\in\mathbb{Z}_{+}$. When $m$ is replaced with $m+1$, we have
$\sum_{k=0}^{m+1}(-1)^{k}\binom{n+1}{k}=\sum_{k=0}^{m}(-1)^{k}\binom{n+1}{k}+(-1)^{m+1}\binom{n+1}{m+1}$ (4.6)
$=$ $(-1)^{m}\binom{n}{m}+(-1)^{m+1}\binom{n+1}{m+1}=(-1)^{m+1}\binom{n}{m+1}$.
The assumption is used in the derivation of the second line. ∎
In the following we denote an expectation value of a random variable $X$ by $\mathbb{E}X$.
Lemma 4.5
Suppose that a $\mathbb{Z}_{+}$ valued random variable $X$ satisfies
$\mathbb{E}\binom{X}{k}=\beta_{k}<\infty$ (4.7)
for arbitrary $k\in\mathbb{Z}_{+}$. Then the following inequality holds for arbitrary $j,\ell\in\mathbb{Z}_{+}$.
$\sum_{k=j}^{j+2\ell+1}(-1)^{k-j}\binom{k}{j}\beta_{k}\leq\mathbb{P}(X=j)\leq\sum_{k=j}^{j+2\ell}(-1)^{k-j}\binom{k}{j}\beta_{k}.$ (4.8)
Proof
Since
$\beta_{k}=\sum_{n=0}^{\infty}\binom{n}{k}\mathbb{P}(X=n),$ (4.9)
we obtain
$\sum_{k=j}^{j+m}(-1)^{k-j}\binom{k}{j}\beta_{k}$ $=$ $\sum_{k=j}^{j+m}(-1)^{k-j}\binom{k}{j}\sum_{n=0}^{\infty}\binom{n}{k}\mathbb{P}(X=n)$ (4.10)
$=$ $\sum_{n=0}^{\infty}\mathbb{P}(X=n)\sum_{k=j}^{j+m}(-1)^{k-j}\binom{k}{j}\binom{n}{k}$
for $j,m\in\mathbb{Z}_{+}$. Using
$\binom{k}{j}\binom{n}{k}=\binom{n}{j}\binom{n-j}{k-j}$ (4.11)
and Lemma 4.4, we further find
$\sum_{k=j}^{j+m}(-1)^{k-j}\binom{k}{j}\beta_{k}=\sum_{n=j}^{\infty}\mathbb{P}(X=n)\binom{n}{j}\sum_{k=j}^{j+m}(-1)^{k-j}\binom{n-j}{k-j}$ (4.12)
$=$ $\mathbb{P}(X=j)+(-1)^{m}\sum_{n=j+1}^{\infty}\mathbb{P}(X=n)\binom{n}{j}\binom{n-j-1}{m}$.
Considering the cases $m$ odd and $m$ even, respectively, we arrive at the required inequality. ∎
Corollary 4.6
Let $\{X_{n}\}_{n\in\mathbb{N}}$ be a sequence of $\mathbb{Z}_{+}$ valued random variables, and for arbitrary $k\in\mathbb{Z}_{+}$ assume that $\lim_{n\rightarrow\infty}\mathbb{E}\binom{X_{n}}{k}$ converges to a finite value $\beta_{k}$. Then, if a series
$\sum_{k=j}^{\infty}(-1)^{k-j}\binom{k}{j}\beta_{k}$ (4.13)
converges, $\lim_{n\rightarrow\infty}\mathbb{P}(X_{n}=j)$ also converges and
$\lim_{n\rightarrow\infty}\mathbb{P}(X_{n}=j)=\sum_{k=j}^{\infty}(-1)^{k-j}\binom{k}{j}\beta_{k}$ (4.14)
holds.
Proof
Let us put $\beta_{k}^{(n)}=\mathbb{E}\binom{X_{n}}{k}$. Since $\lim_{n\rightarrow\infty}\beta_{k}^{(n)}<\infty$, it can be seen that $\beta_{k}^{(n)}<\infty$ for sufficiently large $n$. Therefore, for such $n$,
$\sum_{k=j}^{j+2\ell+1}(-1)^{k-j}\binom{k}{j}\beta_{k}^{(n)}\leq\mathbb{P}(X_{n}=j)\leq\sum_{k=j}^{j+2\ell}(-1)^{k-j}\binom{k}{j}\beta_{k}^{(n)}$ (4.15)
holds due to Lemma 4.5. Taking the limit $n\rightarrow\infty$, we find
$\sum_{k=j}^{j+2\ell+1}(-1)^{k-j}\binom{k}{j}\beta_{k}\leq\liminf_{n\rightarrow\infty}\mathbb{P}(X_{n}=j)$ (4.16)
$\leq$ $\limsup_{n\rightarrow\infty}\mathbb{P}(X_{n}=j)\leq\sum_{k=j}^{j+2\ell}(-1)^{k-j}\binom{k}{j}\beta_{k}.$
We finally take the limit $\ell\rightarrow\infty$ to obtain the required result. ∎
Let us now introduce a formula giving an evaluation of $\mathbb{E}\binom{X}{k}$ for a special $\mathbb{Z}_{+}$ valued random variable $X$.
Lemma 4.7
Suppose that $B_{1},B_{2},\cdots,B_{m}$ are $\{0,1\}$-valued random variables. When a $\mathbb{Z}_{+}$-valued random variable $X$ is defined as
$X=\sum_{j=1}^{m}B_{j},$ (4.17)
an equality
$\mathbb{E}\binom{X}{k}=\sum_{\begin{subarray}{c}I\subset S\\ \#I=k\end{subarray}}\mathbb{E}\left\{\prod_{j\in I}B_{j}\right\}$ (4.18)
holds for $S=\{1,2,\cdots,m\}$ and $k\in\mathbb{N}$.
Proof
Let us first prove an identity
$\prod_{j\in I}B_{j}=\sum_{\begin{subarray}{c}J\subset S\\ J\supset I\end{subarray}}\prod_{j\in J}B_{j}\prod_{\ell\in S\setminus J}(1-B_{\ell})$ (4.19)
for $I\subset S$. If $\prod_{j\in I}B_{j}=0$, the both sides of (4.19) are obviously zero. If $\prod_{j\in I}B_{j}=1$, each term in the sum over $J$ in the right hand side is zero, unless a condition $B_{j}=1$ ($j\in J$) and $B_{\ell}=0$ ($\ell\in S\setminus J$) is satisfied. Only a single term satisfies such a condition and its value is one. Therefore the both sides of (4.19) are one.
Using the identity (4.19), one can see that
$\sum_{\begin{subarray}{c}I\subset S\\ \#I=k\end{subarray}}\mathbb{E}\left\{\prod_{j\in I}B_{j}\right\}$ $=$ $\sum_{\begin{subarray}{c}I\subset S\\ \#I=k\end{subarray}}\sum_{\begin{subarray}{c}J\subset S\\ J\supset I\end{subarray}}\mathbb{E}\left\{\prod_{j\in J}B_{j}\prod_{\ell\in S\setminus J}(1-B_{\ell})\right\}$ (4.20)
$=$ $\sum_{\begin{subarray}{c}J\subset S\\ \#J\geq k\end{subarray}}\sum_{\begin{subarray}{c}I\subset J\\ \#I=k\end{subarray}}\mathbb{E}\left\{\prod_{j\in J}B_{j}\prod_{\ell\in S\setminus J}(1-B_{\ell})\right\}$
$=$ $\sum_{n=k}^{m}\sum_{\begin{subarray}{c}J\subset S\\ \#J=n\end{subarray}}\binom{n}{k}\mathbb{E}\left\{\prod_{j\in J}B_{j}\prod_{\ell\in S\setminus J}(1-B_{\ell})\right\}.$
On the other hand, we can write the probability $\mathbb{P}(X=x)$ as
$\mathbb{P}(X=n)=\sum_{\begin{subarray}{c}J\subset S\\ \#J=n\end{subarray}}\mathbb{E}\left\{\prod_{j\in J}B_{j}\prod_{\ell\in S\setminus J}(1-B_{\ell})\right\}.$ (4.21)
Therefore we find
$\sum_{\begin{subarray}{c}I\subset S\\ \#I=k\end{subarray}}\mathbb{E}\left\{\prod_{j\in I}B_{j}\right\}=\sum_{n=k}^{m}\binom{n}{k}\mathbb{P}(X=n)=\mathbb{E}\binom{X}{k}$, (4.22)
which gives the required result. ∎
A necessary condition for a bipartite graph $G=({\cal V}_{1},{\cal V}_{2},{\cal E})$ with ${\cal V}_{1}=\#{\cal V}_{2}=n$ to have a perfect matching is the absence of isolated points defined below.
Definition 4.8
In a bipartite graph $G=({\cal V}_{1},{\cal V}_{2},{\cal E})$, $v\in{\cal V}_{1}$ satisfying
$(v,w)\notin{\cal E}\ {\rm for}\ {\rm arbitrary}\ w\in{\cal V}_{2}$
and $w\in{\cal V}_{2}$ satisfying
$(v,w)\notin{\cal E}\ {\rm for}\ {\rm arbitrary}\ v\in{\cal V}_{1}$
are called isolated points of $G$.
We first evaluate the probability to have no isolated points, because it will eventually turn out to dominate the probability to have a perfect matching in the limit $n\rightarrow\infty$.
Lemma 4.9
Let us come back to the notations of Theorem 4.2 and again consider a random bipartite graph $G_{n}=({\cal V}_{1},{\cal V}_{2},{\cal E})$ with ${\cal E}\subset{\cal V}_{1}\times{\cal V}_{2}$ and ${\cal V}_{1}=\#{\cal V}_{2}=n$. Each $\langle j,\ell\rangle\in{\cal V}_{1}\times{\cal V}_{2}$ ($j,\ell=1,2,\cdots,n$) independently satisfies $\langle j,\ell\rangle\in{\cal E}$ with probability $p$. The probability parameter $p$ depends on $n$ as
$p=p(n)=\frac{\log n+c}{n}+o\left(\frac{1}{n}\right),\ \ \ n\rightarrow\infty,$ (4.23)
where $c\in\mathbb{R}$ is a constant. Suppose that the number of the isolated points of the random bipartite graph $G_{n}=({\cal V}_{1},{\cal V}_{2},{\cal E})$ is $X_{n}$. The probability to have $X_{n}=x$ for a fixed $x\in\mathbb{Z}_{+}$ is asymptotically evaluated as
$\lim_{n\rightarrow\infty}\mathbb{P}(X_{n}=x)=e^{-\lambda}\frac{\lambda^{x}}{x!}$ (4.24)
with $\lambda=2e^{-c}$.
Proof
It follows from Lemma 4.7 that
$\mathbb{E}\binom{X_{n}}{k}=\sum_{\begin{subarray}{c}I\subset{\cal V}_{1}\cup{\cal V}_{2}\\ \#I=k\end{subarray}}\mathbb{P}\left({\rm the}\ {\rm elements}\ {\rm of}\ I\ {\rm are}\ {\rm all}\ {\rm isolated}\ {\rm points}\ {\rm of}\ G_{n}\right).$ (4.25)
Let us write $I\subset{\cal V}_{1}\cup{\cal V}_{2}$ with #I=k\#I=k as
$I=I_{1}\cup I_{2},\ \ \ I_{1}\subset{\cal V}_{1},\ \ \ I_{2}\subset{\cal V}_{2}$ (4.26)
with
$I_{1}=k_{1}\geq 0,\ \ \ \#I_{2}=k_{2}\geq 0,\ \ \ k_{1}+k_{2}=k.$ (4.27)
There are $\binom{n}{k_{1}}\binom{n}{k_{2}}$ ways to choose such $\{I_{1},I_{2}\}$. Then the probabilities to have $(v,w)\notin{\cal E}$ for arbitrary $(v,w)\in I_{1}\times I_{2}$, for arbitrary $(v,w)\in I_{1}\times\left({\cal V}_{2}\setminus I_{2}\right)$ and for arbitrary $(v,w)\in\left({\cal V}_{1}\setminus I_{1}\right)\times I_{2}$ are
$(1-p)^{k_{1}k_{2}},(1-p)^{k_{1}(n-k_{2})}\ {\rm and}\ (1-p)^{k_{2}(n-k_{1})},$ (4.28)
respectively. Hence we obtain
$\mathbb{E}\binom{X_{n}}{k}$ $=$ $\sum_{k_{1}+k_{2}=k}\binom{n}{k_{1}}\binom{n}{k_{2}}(1-p)^{k_{1}k_{2}}(1-p)^{k_{1}(n-k_{2})}(1-p)^{k_{2}(n-k_{1})}$ (4.29)
$=$ $\sum_{k_{1}+k_{2}=k}\binom{n}{k_{1}}\binom{n}{k_{2}}(1-p)^{nk-k_{1}k_{2}}.$
One can see from (4.23) that
$\log(1-p)=-\frac{\log n+c}{n}+o\left(\frac{1}{n}\right),$ (4.30)
which leads to
$\mathbb{E}\binom{X_{n}}{k}$ $=$ $\sum_{k_{1}+k_{2}=k}\frac{n^{k}(1+o(1))}{k_{1}!k_{2}!}{\rm exp}\left\{(nk-k_{1}k_{2})\log(1-p)\right\}$ (4.31)
$=$ $\sum_{k_{1}+k_{2}=k}\frac{1+o(1)}{k_{1}!k_{2}!}{\rm exp}\left\{k\log n+nk\log(1-p)+o(1)\right\}$
$=$ $\frac{1}{k!}\sum_{k_{1}=0}^{k}\binom{k}{k_{1}}{\rm exp}\left(-ck\right)+o(1)$
$=$ $\frac{2^{k}e^{-ck}}{k!}+o(1)=\frac{\lambda^{k}}{k!}+o(1).$
Now we can utilize Corollary 4.6 to find
$\lim_{n\rightarrow\infty}\mathbb{P}(X_{n}=x)$ $=$ $\sum_{k=x}^{\infty}(-1)^{k-x}\binom{k}{x}\frac{\lambda^{k}}{k!}=\sum_{k=x}^{\infty}\frac{(-1)^{k-x}\lambda^{k}}{(k-x)!x!}$ (4.32)
$=$ $\frac{\lambda^{x}}{x!}\sum_{\ell=0}\frac{(-\lambda)^{\ell}}{\ell!}=\frac{\lambda^{x}}{x!}e^{-\lambda}.$
In order to prove Theorem 4.2, we need to show that the probability for a random bipartite graph $G$ to have neither isolated points nor perfect matchings converges to zero.
When a bipartite graph $G=({\cal V}_{1},{\cal V}_{2},{\cal E})$ has no perfect matchings, Theorem 3.6 (Hall’s marriage theorem) means that
there exists $A_{1}\subset{\cal V}_{1}$ satisfying $\Gamma_{G}(A_{1})\leq\#A_{1}-1$.
The same theorem can also be applied to the case when ${\cal V}_{1}$ and ${\cal V}_{2}$ are interchanged and consequently
there exists $A_{2}\subset{\cal V}_{2}$ satisfying $\Gamma_{G}(A_{2})\leq\#A_{2}-1$.
As for such sets $A_{1}$ and $A_{2}$, we have the following lemma.
Lemma 4.10
Let us consider a bipartite graph $G=({\cal V}_{1},{\cal V}_{2},{\cal E})$ with ${\cal V}_{1}={\cal V}_{2}=n$ and suppose that $G$ has no perfect matchings. Among the sets $A\subset{\cal V}_{1}$ and $A\subset{\cal V}_{2}$ which satisfy $\Gamma_{G}(A)\leq\#A-1$, we choose one set minimizing #A\#A and call it $I$. Then $I$ satisfies the followings.
(1) $\Gamma_{G}(I)=\#I-1$.
(2) $I\leq(n+1)/2$.
(3) $(\Gamma_{G}(w)\cap I)\geq 2$ for arbitrary $w\in\Gamma_{G}(I)$.
Proof
(1) We arbitrarily choose $v\in I$ and put $I^{\prime}=I\setminus\{v\}$. As $I$ minimizes the number of elements in the family of sets unsatisfying Hall’s condition, one finds $\Gamma_{G}(I^{\prime})\geq\#I^{\prime}$. Moreover, as $I^{\prime}\subset I$, we obtain $\Gamma_{G}(I^{\prime})\subset\Gamma_{G}(I)$. It follows that
$\Gamma_{G}(I)\geq\#\Gamma_{G}(I^{\prime})\geq\#I^{\prime}=\#I-1.$ (4.33)
On the other hand, since $I$ does not satisfy Hall’s condition, one sees $\Gamma_{G}(I)\leq\#I-1$. Therefore the required result together with the equality $\Gamma_{G}(I)=\Gamma_{G}(I^{\prime})$ holds.
(2) Let us first assume $I\subset{\cal V}_{1}$. If #I>(n+1)/2\#I>(n+1)/2, then $\Gamma_{G}(I)=\#I-1>(n-1)/2$ from (1). We now put $K={\cal V}_{2}\setminus\Gamma_{G}(I)$ and find $K=n-\#\Gamma_{G}(I)<(n+1)/2$. On the other hand, as $\Gamma_{G}(K)\subset{\cal V}_{1}\setminus I$,
$\Gamma_{G}(K)\leq n-\#I=n-\#\Gamma_{G}(I)-1=\#K-1.$ (4.34)
This is contradictory because $K$ does not satisfy Hall’s condition in spite of the fact that #K<(n+1)/2<#I\#K<(n+1)/2<\#I. A similar argument holds in the other case $I\subset{\cal V}_{2}$.
(3) We arbitrarily choose $w\in\Gamma_{G}(I)$. From the definition of $\Gamma_{G}$, there exists $v\in I$ and $(v,w)\in{\cal E}$. Now we define $I^{\prime}=I\setminus\{v\}$, then $\Gamma_{G}(I)=\Gamma_{G}(I^{\prime})$ from (1). Hence $v^{\prime}\in I^{\prime}$ exists, so that $(v^{\prime},w)\in{\cal E}$. Therefore $(\Gamma_{G}(w)\cap I)\geq 2$. ∎
Due to Lemma 4.10, if a bipartite graph $G$ has no perfect matchings, then it has at least one isolated point or satisfies the following condition.
Condition 4.11
There exists $k\in\mathbb{N}$ ($2\leq k\leq(n+1)/2$) satisfying $I\subset{\cal V}_{1}$, $J\subset{\cal V}_{2}$ (or $I\subset{\cal V}_{2}$, $J\subset{\cal V}_{1}$) with #I=k\#I=k, #J=k1\#J=k-1, so that $\Gamma_{G}(I)=J$ and $\forall w\in J$, $(\Gamma_{G}(w)\cap I)\geq 2$.
The case $k=1$ corresponds to graphs with at least one isolated point. Now we asymptotically evaluate the probability that Condition 4.11 holds.
Lemma 4.12
Let us suppose ${\cal V}_{1}=\#{\cal V}_{2}=n$ and consider a random bipartite graph $G_{n}({\cal V}_{1},{\cal V}_{2},{\cal E})$. Each $\langle j,\ell\rangle\in{\cal V}_{1}\times{\cal V}_{2}$ ($j,\ell=1,2,\cdots,n$) independently satisfies $\langle j,\ell\rangle\in{\cal E}$ with probability $p$. The probability parameter $p$ depends on $n$ as
$p=p(n)=\frac{\log n+c}{n}+o\left(\frac{1}{n}\right),\ \ \ n\rightarrow\infty$ (4.35)
with a constant $c\in\mathbb{R}$. Then
$\lim_{n\rightarrow\infty}\mathbb{P}\left(G_{n}\ {\rm satisfies}\ {\bf Condition\ 4.11}\right)=0$ (4.36)
holds.
Proof
Let us first fix $k\in\mathbb{N}$ satisfying $2\leq k\leq(n+1)/2$. We choose vertex sets $I\subset{\cal V}_{1}$ and $J\subset{\cal V}_{2}$ with #I=k\#I=k and #J=k1\#J=k-1. There are $\binom{n}{k}\binom{n}{k-1}$ ways to choose such $\{I,J\}$. Then the probability to have $(v,w)\notin{\cal E}$ for all $(v,w)\in I\times{\cal V}_{2}\setminus J$ is $(1-p)^{k(n-k+1)}$. It follows that
$\mathbb{P}(\Gamma_{G}(I)=J)\leq(1-p)^{k(n-k+1)}$ (4.37)
for each $\{I,J\}$.
Next we fix a vertex $w\in J$ and choose two distinct vertices $v,v^{\prime}\in I$. For each $w$ there are $\binom{k}{2}$ ways to choose such $v,v^{\prime}$. The probability to have $(v,w)\in{\cal E}$ and $(v^{\prime},w)\in{\cal E}$ for each triple $\{w,v,v^{\prime}\}$ is $p^{2}$. Thus we obtain
$\mathbb{P}(\exists\{v,v^{\prime}\},(v,w)\in{\cal E}\ {\rm and}\ (v^{\prime},w)\in{\cal E})\leq\binom{k}{2}p^{2}$ (4.38)
for each $w$.
We can see from the above argument that
$\mathbb{P}\left(G_{n}\ {\rm satisfies}\ {\bf Condition\ 4.11}\right)$ (4.39)
$\leq$ $2\sum_{k=2}^{\lfloor(n+1)/2\rfloor}\binom{n}{k}\binom{n}{k-1}\left\{\binom{k}{2}p^{2}\right\}^{k-1}(1-p)^{k(n-k+1)}$
$\leq$ $2\sum_{k=2}^{\lfloor(n+1)/2\rfloor}\left(\frac{ne}{k}\right)^{k}\left(\frac{ne}{k-1}\right)^{k-1}\left\{\binom{k}{2}p^{2}\right\}^{k-1}e^{-pk(n-k+1)}.$
Note that there is a factor $2$ in front of the sum over $k$ due to another case $I\in{\cal V}_{2}$ and $J\in{\cal V}_{1}$. The first term with $k=2$ is
$\frac{(ne)^{3}}{2}p^{2}e^{-2p(n-1)}$ (4.40)
$=$ $\frac{(ne)^{3}}{2}\left\{\frac{\log n+c}{n}+o\left(\frac{1}{n}\right)\right\}^{2}{\rm exp}\left[-2\left\{\frac{\log n+c}{n}+o\left(\frac{1}{n}\right)\right\}(n-1)\right]$
$=$ $O\left(\frac{(\log n)^{2}}{n}\right)\rightarrow 0,\ \ \ n\rightarrow\infty.$
The sum of the other terms with $k\geq 3$ is
$\varOmega$ $=$ $2\sum_{k=3}^{\lfloor(n+1)/2\rfloor}\left(\frac{ne}{k}\right)^{k}\left(\frac{ne}{k-1}\right)^{k-1}\left\{\binom{k}{2}p^{2}\right\}^{k-1}e^{-pk(n-k+1)}$ (4.41)
$=$ $\frac{4}{nep^{2}}\sum_{k=3}^{\lfloor(n+1)/2\rfloor}\frac{1}{k}\left(\frac{(nep)^{2}}{2}\right)^{k}e^{-pk(n-k+1)}.$
Here one can see from $k\leq(n+1)/2$ that
$e^{-p(n-k+1)}\leq e^{-p(n+1)/2}\leq e^{-pn/2}.$ (4.42)
Thus we find
$\varOmega\leq\frac{1}{np^{2}}\sum_{k=3}^{\lfloor(n+1)/2\rfloor}u^{k}\leq\frac{1}{np^{2}}\frac{u^{3}}{1-u},$ (4.43)
where
$u=\frac{(nep)^{2}}{2}e^{-pn/2}.$ (4.44)
Let us take constants $c_{1},c_{2}\in\mathbb{R}$ satisfying
$\frac{\log n+c_{1}}{n}\leq p(n)\leq\frac{\log n+c_{2}}{n}$ (4.45)
for sufficiently large $n$. Then we obtain
$\frac{1}{np^{2}}\leq\frac{n}{(\log n+c_{1})^{2}}$ (4.46)
and
$u\leq\frac{n^{2}e^{2}}{2}\left(\frac{\log n+c_{2}}{n}\right)^{2}{\rm exp}\left\{-\frac{\log n+c_{1}}{2}\right\}=\frac{e^{2-c_{1}/2}}{2}\frac{(\log n+c_{2})^{2}}{\sqrt{n}}.$ (4.47)
It follows from (4.43), (4.46) and (4.47) that
$\varOmega$ $\leq$ $\frac{n}{(\log n+c_{1})^{2}}\frac{e^{6-3c_{1}/2}}{8}\frac{(\log n+c_{2})^{6}}{n\sqrt{n}}\left\{1-\frac{e^{2-c_{1}/2}}{2}\frac{(\log n+c_{2})^{2}}{\sqrt{n}}\right\}^{-1}$ (4.48)
$=$ $\frac{(\log n+c_{2})^{6}}{\sqrt{n}(\log n+c_{1})^{2}}O(1)\rightarrow 0,\ \ \ n\rightarrow\infty.$
Proof of Theorem 4.2
As discussed above, for a bipartite graph $G_{n}=({\cal V}_{1},{\cal V}_{2},{\cal E})$ with $V_{1}=\#V_{2}=n$,
$\mathbb{P}(G_{n}\ {\rm has}\ {\rm at}\ {\rm least}\ {\rm one}\ {\rm isolated}\ {\rm point})\leq\mathbb{P}(G_{n}\ {\rm has}\ {\rm no}\ {\rm perfect}\ {\rm matchings})$
$\leq$ $\mathbb{P}(G_{n}\ {\rm has}\ {\rm at}\ {\rm least}\ {\rm one}\ {\rm isolated}\ {\rm point})+\mathbb{P}\left(G_{n}\ {\rm satisfies}\ {\bf Condition\ 4.11}\right)$
holds. We can see from Lemma 4.9, Lemma 4.12 and the squeeze theorem that
$\lim_{n\rightarrow\infty}\mathbb{P}(G_{n}\ {\rm has}\ {\rm no}\ {\rm perfect}\ {\rm matchings})$ (4.50)
$=$ $\lim_{n\rightarrow\infty}\mathbb{P}(G_{n}\ {\rm has}\ {\rm at}\ {\rm least}\ {\rm one}\ {\rm isolated}\ {\rm point})$
$=$ $1-e^{-\lambda}.$
4.2 Eigenvalue degeneracy probability for sparse random matrices
We are now in a position to begin the proof of the main theorem. Suppose that random variables $Y_{j\ell}$ ($j,\ell=1,2,\cdots,N)$ independently obey Bernoulli distribution with a parameter $p$ and $K$-valued (broad) continuous random variables $Z_{j\ell}$ ($j,\ell=1,2,\cdots,N$) are independently distributed. We want to evaluate the degeneracy probability for the eigenvalues of the $N\times N$ random matrix $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$.
Theorem 4.13
Let $c\in\mathbb{R}$. A parameter $p$ depends on $N$ and satisfies
$p=p(N)=\frac{\log N+c}{N}+o\left(\frac{1}{N}\right),\ \ \ N\rightarrow\infty.$ (4.51)
Then
$\lim_{N\rightarrow\infty}\mathbb{P}\left({\rm the}\ {\rm matrix}\ (Y_{jl}Z_{jl})\ {\rm has}\ N\ {\rm distinct}\ {\rm eigenvalues}\right)=e^{-\lambda}+\lambda e^{-\lambda}$ (4.52)
holds, where $\lambda=2e^{-c}$.
Proof
We evaluate the probability for the corresponding random bipartite graph $G=({\cal V}^{R},{\cal V}^{C},{\cal E})$ with ${\cal E}=\{(j,\ell)|Y_{j\ell}=1\}$ to satisfy Condition 4.1, because it is equal to the probability that $N$ eigenvalues are all distinct.
Let us first bound the probability above. We write the number of isolated points of $G$ as $X_{N}$ and $\{1,2,\cdots,N\}$ as ${\cal I}_{N}$. Then we find
$\mathbb{P}(G\ {\rm satisfies}\ {\bf Condition\ 4.1})$ (4.53)
$=$ $\mathbb{P}(G\ {\rm has}\ {\rm a}\ {\rm perfect}\ {\rm matching}$
${\rm or}\ \exists j\in{\cal I}_{N},\ G[{\cal I}_{N}\setminus\{j\}]\ {\rm has}\ {\rm a}\ {\rm perfect}\ {\rm matching}).$
As a necessary condition to have a perfect matching is the absence of isolated points, one can see that
$\mathbb{P}(G\ {\rm satisfies}\ {\bf Condition\ 4.1})$
$\leq$ $\mathbb{P}(G\ {\rm has}\ {\rm no}\ {\rm isolated}\ {\rm points}$
${\rm or}\ \exists j\in{\cal I}_{N},G[{\cal I}_{N}\setminus\{j\}]\ {\rm has}\ {\rm no}\ {\rm isolated}\ {\rm points})$
$=$ $\mathbb{P}(X_{N}=0)$
$+\mathbb{P}(X_{N}\geq 1\ {\rm and}\ \exists j\in{\cal I}_{N},G[{\cal I}_{N}\setminus\{j\}]\ {\rm has}\ {\rm no}\ {\rm isolated}\ {\rm points}).$
When $X_{N}\geq 1$ and $\exists j\in{\cal I}_{N}$, $G[{\cal I}_{N}\setminus\{j\}]$ has no isolated points, the isolated points of $G$ must be chosen from $\{v^{R}_{j},v^{C}_{j}\}$. It follows that
$\mathbb{P}(G\ {\rm satisfies}\ {\bf Condition\ 4.1})$
$\leq$ $\mathbb{P}(X_{N}\leq 1)+\mathbb{P}\left(\exists j\in{\cal I}_{N},v^{R}_{j}\ {\rm and}\ v^{C}_{j}\ {\rm are}\ {\rm both}\ {\rm isolated}\ {\rm points}\right).$
Because of Lemma 4.9, the first term of the RHS satisfies
$\lim_{N\rightarrow\infty}\mathbb{P}(X_{N}\leq 1)=e^{-\lambda}+\lambda e^{-\lambda}.$ (4.56)
The second term of the RHS is estimated as
$\mathbb{P}\left(\exists j\in{\cal I}_{N},v^{R}_{j}\ {\rm and}\ v^{C}_{j}\ {\rm are}\ {\rm both}\ {\rm isolated}\ {\rm points}\right)$ (4.57)
$\leq$ $\sum_{j\in{\cal I}_{N}}\mathbb{P}\left(v^{R}_{j}\ {\rm and}\ v^{C}_{j}\ {\rm are}\ {\rm both}\ {\rm isolated}\ {\rm points}\right)$
$=$ $\sum_{j\in{\cal I}_{N}}(1-p)^{2N-1}=N(1-p)^{2N-1}.$
Then one can see that
$\mathbb{P}\left(\exists j\in{\cal I}_{N},v^{R}_{j}\ {\rm and}\ v^{C}_{j}\ {\rm are}\ {\rm both}\ {\rm isolated}\ {\rm points}\right)$ (4.58)
$\leq$ ${\rm exp}\left\{\log N+(2N-1)\log(1-p)\right\}$
$=$ ${\rm exp}\left[\log N-(2N-1)\left\{\frac{\log N+c}{N}+o\left(\frac{1}{N}\right)\right\}\right]$
$=$ ${\rm exp}\left\{-\log N+O(1)\right\}\rightarrow 0,\ \ \ N\rightarrow\infty.$
Thus we obtain
$\lim_{N\rightarrow\infty}\mathbb{P}(G\ {\rm satisfies}\ {\bf Condition\ 4.1})\leq e^{-\lambda}+\lambda e^{-\lambda}.$ (4.59)
Next we need to bound the probability below as
$\lim_{N\rightarrow\infty}\mathbb{P}(G\ {\rm satisfies}\ {\bf Condition\ 4.1})\geq e^{-\lambda}+\lambda e^{-\lambda},$ (4.60)
or equivalently
$\lim_{N\rightarrow\infty}\mathbb{P}(G\ {\rm does}\ {\rm not}\ {\rm satisfy}\ {\bf Condition\ 4.1})\leq 1-e^{-\lambda}-\lambda e^{-\lambda}.$ (4.61)
In order to prove (4.61), we classify the cases according to the number of isolated points.
(1) the case $X_{N}\geq 2$
We can see from Lemma 4.9 that the corresponding asymptotic probability is
$\lim_{N\rightarrow\infty}\mathbb{P}(X_{N}\geq 2)=1-e^{-\lambda}-\lambda e^{-\lambda}.$ (4.62)
Therefore we obtain
$\lim_{N\rightarrow\infty}\mathbb{P}(X_{N}\geq 2\ {\rm and}\ G\ {\rm does}\ {\rm not}\ {\rm satisfy}\ {\bf Condition\ 4.1})$ (4.63)
$\leq$ $\lim_{N\rightarrow\infty}\mathbb{P}(X_{N}\geq 2)=1-e^{-\lambda}-\lambda e^{-\lambda}.$
(2) the case $X_{N}=1$
Suppose that $v^{R}_{k}\in{\cal V}_{R}$ is the only isolated point of $G$. In order to make Condition 4.1 unsatisfied, a bipartite subgraph $G[{\cal I}_{N}\setminus\{k\}]$ must have no perfect matchings. Then $G[{\cal I}_{N}\setminus\{k\}]$ has at least an isolated point or Condition 4.11 holds. Let us separately consider those two cases.
(a) If $G[{\cal I}_{N}\setminus\{k\}]$ has an isolated point $w$, then $w$ is not an isolated point of $G$. Therefore $w\in{\cal V}_{R}$ and $(w,v^{C}_{N})\in{\cal E}$. It follows that the total probability to have such pairs $\{v^{R}_{k},w\}$ is equal to or less than
$\Pi_{a}(N)=\sum_{k=1}^{N}\mathbb{P}(v^{R}_{k}\ {\rm is}\ {\rm an}\ {\rm isolated}\ {\rm point}\ {\rm of}\ G)$
$\times\sum_{j\in{\cal I}_{N}\setminus\{k\}}\mathbb{P}((v^{R}_{j},v^{C}_{k})\in{\cal E})\ \mathbb{P}(v^{R}_{j}\ {\rm is}\ {\rm an}\ {\rm isolated}\ {\rm point}\ {\rm of}\ G[{\cal I}_{N}\setminus\{k\}]).$
We can see from
$\mathbb{P}((v^{R}_{j},v^{C}_{k})\in{\cal E})=p(N)$ (4.65)
and
$\mathbb{P}(v^{R}_{k}\ {\rm is}\ {\rm an}\ {\rm isolated}\ {\rm point}\ {\rm of}\ G)=(1-p)^{N}=\frac{\mathbb{E}[X_{N}]}{2N}$ (4.66)
that
$\Pi_{a}(N)=\frac{1}{4}p(N)\mathbb{E}[X_{N}]\mathbb{E}[X_{N-1}].$ (4.67)
In the proof of Lemma 4.9, we showed that $\mathbb{E}[X_{N}]<\infty$. Moreover, as
$p(N)=\frac{\log N+c}{N}+o\left(\frac{1}{N}\right)=p(N-1)+o\left(\frac{1}{N}\right),$ (4.68)
$\mathbb{E}[X_{N-1}]<\infty$ also holds. Therefore $\Pi_{a}(N)$ goes to zero in the limit $N\rightarrow\infty$. We need to consider the other case when $v^{C}_{k}\in{\cal V}_{C}$ is the only isolated point of $G$. In that case, a similar argument can be applied and we find that the corresponding probability also goes to zero in the limit $N\rightarrow\infty$.
(b) Next we consider the case when $v^{R}_{k}\in{\cal V}_{R}$ is the only isolated point of $G$ and $G[{\cal I}_{N}\setminus\{k\}]$ satisfies Condition 4.11. The total probability of such cases is equal to or less than
$\Pi_{b}(N)$ $=$ $\sum_{k=1}^{N}\mathbb{P}(v^{R}_{k}\ {\rm is}\ {\rm an}\ {\rm isolated}\ {\rm point}\ {\rm of}\ G)$ (4.69)
$\times\mathbb{P}(G[{\cal I}_{N}\setminus\{k\}]\ {\rm satisfies}\ {\bf Condition\ 4.11})$
$=$ $\frac{1}{2}\mathbb{E}[X_{N}]\ \mathbb{P}(G[{\cal I}_{N}\setminus\{k\}]\ {\rm satisfies}\ {\bf Condition\ 4.11}).$
We can again see from the proof of Lemma 4.9 that $\mathbb{E}[X_{N}]<\infty$. Since p(N)=p(N1)+o(1/N)p(N)=p(N-1)+o(1/N), Lemma 4.12 leads to
$\lim_{N\rightarrow\infty}\mathbb{P}(G[{\cal I}_{N}\setminus\{k\}]\ {\rm satisfies}\ {\bf Condition\ 4.11})=0.$ (4.70)
Thus $\Pi_{b}(N)$ goes to zero in the limit $N\rightarrow\infty$. The other case when $v^{C}_{k}\in{\cal V}_{C}$ can similarly be treated and we find that the corresponding probability also goes to zero in the limit $N\rightarrow\infty$.
Putting the estimates in the cases (a) an (b) together, we find
$\lim_{N\rightarrow\infty}\mathbb{P}(X_{N}=1\ {\rm and}\ G\ {\rm does}\ {\rm not}\ {\rm satisfy}\ {\bf Condition\ 4.1})$ (4.71)
$\leq$ $\lim_{N\rightarrow\infty}(2\Pi_{a}(N)+2\Pi_{b}(N))=0.$
(3) the case $X_{N}=0$
If $X_{N}=0$, in order to make Condition 4.1 unsatisfied, Condition 4.11 must be satisfied. It follows from Lemma 4.12 that
$\lim_{N\rightarrow\infty}\mathbb{P}(G\ {\rm satisfies}\ {\bf Condition\ 4.11})=0.$ (4.72)
Hence one can see that
$\lim_{N\rightarrow\infty}\mathbb{P}(X_{N}=0\ {\rm and}\ G\ {\rm does}\ {\rm not}\ {\rm satisfy}\ {\bf Condition\ 4.1})$ (4.73)
$\leq$ $\lim_{N\rightarrow\infty}\mathbb{P}(G\ {\rm satisfies}\ {\bf Condition\ 4.11})=0.$
It follows from (4.63), (4.71) and (4.73) that (4.61) holds. Therefore we arrive at
$\lim_{N\rightarrow\infty}\mathbb{P}(G\ {\rm satisfies}\ {\bf Condition\ 4.1})=e^{-\lambda}+\lambda e^{-\lambda}.$ (4.74)
This is the required result. ∎
The dependence of the parameter $p$ on $N$ means that the random matrix is sparse. This theorem assures that the eigenvalue degeneracy probability can be positive in such a sparse limit.
5 Eigenvalue degeneracy in symmetric sparse random matrices
In this section, we consider a restricted case when the random matrices are symmetric. Let us put $K=\mathbb{R}$ or $K=\mathbb{C}$. We suppose that $Z_{j\ell}$ ($1\leq j\leq\ell\leq N$) are independent $K$-valued (broad) continuous random variables. Moreover we define $Y_{j\ell}$ ($1\leq j\leq\ell\leq N$) as independent random variables obeying Bernoulli distribution with
$\mathbb{P}(Y_{j\ell}=1)=\left\{\begin{array}[]{ll}p,&{\rm if}\ j<\ell,\\ q,&{\rm if}\ j=\ell.\end{array}\right$. (5.1)
For $1\leq\ell<j\leq N$, we assume $Z_{j\ell}=Z_{\ell j}$ and $Y_{j\ell}=Y_{\ell j}$. Using these random variables, we consider the probability for a symmetric random matrix $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$ to have $N$ distinct eigenvalues. As is well-known, in the case $K=\mathbb{R}$, all the eigenvalues are real.
To begin with, we arbitrarily fix $y_{j\ell}\in\{0,1\}$ ($1\leq j\leq\ell\leq N$) and study the conditional probability under the condition $Y_{j\ell}=y_{j\ell}$. It follows from Theorem 2.5 that this conditional probability is zero or one, and the necessary and sufficient condition to have the conditional probability one is:
There exists at least one set $\{z_{j\ell}\}_{1\leq j\leq\ell\leq N}$ with $z_{j\ell}\in K$, so that the matrix $(y_{j\ell}z_{j\ell})_{j\ell=1,2,\cdots,N}$ has $N$ distinct eigenvalues. Here we assume $z_{j\ell}=z_{\ell j}$ and $y_{j\ell}=y_{\ell j}$ for $1\leq\ell<j\leq N$.
Therefore, as before, the probability for the matrix $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$ to have $N$ distinct eigenvalues does not depend on the distributions of $Z_{j\ell}$ and depends only on $N$, $p$ and $q$. Moreover, from Theorem 3.11, one can see that such a probability is equal to the probability for the random bipartite graph $G=({\cal V}^{R},{\cal V}^{C},\mathfrak{E})$, $\mathfrak{E}=\{\langle j,\ell\rangle|Y_{j\ell}=1\}$ to satisfy Condition 4.1, when ${\cal E}=\mathfrak{E}$.
Let us now find the number distribution of the isolated points. Since the edges in $\mathfrak{E}$ satisfy
$\langle j,\ell\rangle\in\mathfrak{E}\Longleftrightarrow\langle\ell,j\rangle\in\mathfrak{E},$ (5.2)
we obtain
$v_{j}^{R}\ {\rm is}\ {\rm an}\ {\rm isolated}\ {\rm point}\Longleftrightarrow v_{j}^{C}\ {\rm is}\ {\rm an}\ {\rm isolated}\ {\rm point}.$ (5.3)
Therefore, using the maps
$j\in{\cal I}_{N}\mapsto v^{R}_{j}\in{\cal V}^{R}\ {\rm and}\ j\in{\cal I}_{N}\mapsto v^{C}_{j}\in{\cal V}^{C},$ (5.4)
we identify ${\cal V}^{R}$, ${\cal V}^{C}$ and ${\cal I}_{N}=\{1,2,\cdots,N\}$. Accordingly we say $j\in{\cal I}_{N}$ is an isolated point, when $v^{R}_{j}$ and $v^{C}_{j}$ are isolated points. Defining a random variable $X$ as
$X=\#\{j\in{\cal I}_{N}|\ j\ {\rm is}\ {\rm an}\ {\rm isolated}\ {\rm point}\},$ (5.5)
we find the following lemma.
Lemma 5.1
Let $c\in\mathbb{R}$ be a constant. The probability parameters $p$ and $q$ depend on $N$ as
$p=p(N)=\frac{\log N+c}{N}+o\left(\frac{1}{N}\right),\ \ \ N\rightarrow\infty,$ (5.6)
$\lim_{N\rightarrow\infty}q(N)=q_{\infty}\in[0,1].$ (5.7)
Then the probability to have $X=x$ is asymptotically evaluated as
$\lim_{n\rightarrow\infty}\mathbb{P}(X=x)=e^{-\mu}\frac{\mu^{x}}{x!}$ (5.8)
with $\mu=(1-q_{\infty})e^{-c}$.
Proof
Let us define random variables
$B_{j}=\left\{\begin{array}[]{ll}1,&{\rm if}\ j\ {\rm is}\ {\rm an}\ {\rm isolated}\ {\rm point},\\ 0,&{\rm otherwise}\end{array}\right$. (5.9)
($j=1,2,\cdots,N$) and choose a subset $I\subset{\cal I}_{N}$ with $I=k\in\mathbb{N}$. Then
$\mathbb{E}\left(\prod_{j\in I}B_{j}\right)$ $=$ $\mathbb{P}(j\in I\ {\rm are}\ {\rm all}\ {\rm isolated}\ {\rm points})$ (5.10)
$=$ $\prod_{h=1}^{k}\left\{(1-q)(1-p)^{k-h}(1-p)^{N-k}\right\}$
$=$ $(1-q)^{k}(1-p)^{k(k-1)/2}(1-p)^{k(N-k)}.$
It follows from Lemma 4.6 that
$\mathbb{E}\binom{X}{k}$ $=$ $\sum_{\begin{subarray}{c}I\subset{\cal I}_{N}\\ \#I=k\end{subarray}}\mathbb{E}\left(\prod_{j\in I}B_{j}\right)$ (5.11)
$=$ $\binom{N}{k}(1-q)^{k}(1-p)^{k(k-1)/2}(1-p)^{k(N-k)}.$
In the limit $N\rightarrow\infty$ with $k\in\mathbb{N}$ fixed, we find
$\mathbb{E}\binom{X}{k}$ $=$ $(1-q)^{k}\frac{N^{k}\left\{1+o(1)\right\}}{k!}{\rm exp}\left[\left\{kN+O(1)\right\}\log(1-p)\right]$ (5.12)
$=$ $(1-q)^{k}\frac{N^{k}\left\{1+o(1)\right\}}{k!}{\rm exp}\left\{-k\log N-ck+o(1)\right\}$
$=$ $\frac{(1-q_{\infty})^{k}}{k!}{\rm exp}(-ck)+o(1)=\frac{\mu^{k}}{k!}+o(1),$
which also holds for $k=0$. Thus one can see from Corollary 4.5 that
$\lim_{N\rightarrow\infty}\mathbb{P}(X=x)$ $=$ $\sum_{j=x}^{\infty}(-1)^{j-x}\binom{j}{x}\frac{\mu^{j}}{j!}=\sum_{j=x}^{\infty}(-1)^{j-x}\frac{\mu^{j}}{x!(j-x)!}$ (5.13)
$=$ $\frac{\mu^{x}}{x!}\sum_{\ell=0}^{\infty}\frac{(-\mu)^{\ell}}{\ell!}=\frac{\mu^{x}}{x!}e^{-\mu}.$
In the case $q_{\infty}=0$, Lemma 5.1 is reduced to a known result[10] (Theorem 3.1). As before, we now need to consider the consequence of the absence of perfect matchings. For that purpose, we first show the following lemma.
Lemma 5.2
Let us suppose that a bipartite graph $G=({\cal V}_{1},{\cal V}_{2},\mathfrak{E})$ with ${\cal V}_{1}=\#{\cal V}_{2}=n$ satisfies
$\langle j,\ell\rangle\in\mathfrak{E}\Longleftrightarrow\langle\ell,j\rangle\in\mathfrak{E}$ (5.14)
and $G$ has no perfect matchings. We introduce a notation
${\tilde{\Gamma}}_{G}(A)=\{\ell\in{\cal I}_{N}|\exists j\in A,\langle j,\ell\rangle\in\mathfrak{E}\}$ (5.15)
for $A\subset{\cal I}_{N}$. As before, if $A$ is a one-point set $\{j\}$, then ${\tilde{\Gamma}}_{G}(\{j\})$ is simply written as ${\tilde{\Gamma}}_{G}(j)$. Among the sets $A\subset{\cal I}_{N}$ satisfying ${\tilde{\Gamma}}_{G}(A)\leq\#A-1$, we choose one set minimizing #A\#A and call it $I$. Then $I$ satisfies the followings.
(1) ${\tilde{\Gamma}}_{G}(I)=\#I-1$.
(2) $I\leq(n+1)/2$.
(3) $({\tilde{\Gamma}}_{G}(\ell)\cap I)\geq 2$ for arbitrary $\ell\in{\tilde{\Gamma}}_{G}(I)$.
(4) $I\cap{\tilde{\Gamma}}_{G}(I)=\emptyset$.
Proof
We can prove (1), (2) and (3) in the same way as Lemma 4.10. A proof of (4) is given below.
Let us define $S=I\cap{\tilde{\Gamma}}_{G}(I)$ and $J=I\setminus S$. When we arbitrarily take $j\in J$ and $s\in S$, we find $j\notin{\tilde{\Gamma}}_{G}(I)$ and $s\in I$. Therefore $\langle s,j\rangle\notin\mathfrak{E}$, so that $\langle j,s\rangle\notin\mathfrak{E}$ or ${\tilde{\Gamma}}_{G}(J)\cap S=\emptyset$.
On the other hand, it follows from $J\subset I$ that ${\tilde{\Gamma}}_{G}(J)\subset{\tilde{\Gamma}}_{G}(I)$. Using this inclusion relation with ${\tilde{\Gamma}}_{G}(J)\cap S=\emptyset$, we find ${\tilde{\Gamma}}_{G}(J)\subset{\tilde{\Gamma}}_{G}(I)\setminus S$. Noting $S=I\cap{\tilde{\Gamma}}_{G}(I)$ and (1), we obtain
$\#{\tilde{\Gamma}}_{G}(J)$ $\leq$ $\#({\tilde{\Gamma}}_{G}(I)\setminus S)=\#{\tilde{\Gamma}}_{G}(I)-\#S$ (5.16)
$=$ $\#I-1-\#S=\#(I\setminus S)-1=\#J-1.$
Among the sets $A\subset{\cal I}_{N}$ satisfying ${\tilde{\Gamma}}_{G}(A)\leq\#A-1$, $I$ is chosen as the set minimizing #A\#A. Therefore we need to have $I\leq\#J$. Then we find $I=J$ or $S=\emptyset$, because $J\subset I$. ∎
As a consequence of Lemma 5.2, if a bipartite graph $G$ has no perfect matchings, it has at least one isolated point or satisfies the following condition.
Condition 5.3
There exists $k\in\mathbb{N}$ ($2\leq k\leq(n+1)/2$) satisfying $I\subset{\cal I}_{n}$, $J\subset{\cal I}_{n}\setminus I$ with #I=k\#I=k, #J=k1\#J=k-1, so that ${\tilde{\Gamma}}_{G}(I)=J$ and $\forall\ell\in J$, $({\tilde{\Gamma}}_{G}(\ell)\cap I)\geq 2$.
The case $k=1$ again corresponds to graphs with at least one isolated point. The probability for Condition 5.3 to hold is asymptotically evaluated as follows.
Lemma 5.4
Suppose that ${\cal V}_{1}=\#{\cal V}_{2}=n$ and the maps
$v_{1}:{\cal I}_{n}\mapsto{\cal V}_{1}\ {\rm and}\ v_{2}:{\cal I}_{n}\mapsto{\cal V}_{2}$ (5.17)
are bijections. Each $(j,\ell)$ ($1\leq j<\ell\leq n$) independently satisfies $(v_{1}(j),v_{2}(\ell))\in\mathfrak{E}\subset{\cal V}_{1}\times{\cal V}_{2}$ with probability $p$. For $(j,\ell)$ ($1\leq\ell<j\leq n$), we set $(v_{1}(j),v_{2}(\ell))\in\mathfrak{E}\Longleftrightarrow(v_{1}(\ell),v_{2}(j))\in\mathfrak{E}$. The probability parameter $p$ depends on $n$ as
$p=p(n)=\frac{\log n+c}{n}+o\left(\frac{1}{n}\right),\ \ \ n\rightarrow\infty$ (5.18)
with a constant $c\in\mathbb{R}$. Then the random bipartite graph $G_{n}=({\cal V}_{1},{\cal V}_{2},\mathfrak{E})$ fulfills
$\lim_{n\rightarrow\infty}\mathbb{P}\left(G_{n}\ {\rm satisfies}\ {\bf Condition\ 5.3}\right)=0.$ (5.19)
Remark
In Lemma 5.4, the probabilities to have $(v_{1}(j),v_{2}(j))\in\mathfrak{E}$ for $1\leq j\leq n$ are arbitrary. Even independence is not required for them.
Proof of Lemma 5.4
Let us first fix $k\in\mathbb{N}$ satisfying $2\leq k\leq(n+1)/2$. We choose vertex sets $I\subset{\cal I}_{n}$ and $J\subset{\cal I}_{n}\setminus I$ with #I=k\#I=k and #J=k1\#J=k-1. There are $\binom{n}{k}\binom{n-k}{k-1}$ ways to choose such $\{I,J\}$. Then the probability to have $\langle j,\ell\rangle\notin\mathfrak{E}$ for all $(j,\ell)\in I\times\left({\cal I}_{n}\setminus J\right)$ is
$\prod_{h=1}^{k}(1-p)^{k-h}(1-p)^{n-2k+1}=(1-p)^{k(k-1)/2+k(n-2k+1)},$ (5.20)
so that
$\mathbb{P}({\tilde{\Gamma}}_{G}(I)=J)\leq(1-p)^{k(k-1)/2+k(n-2k+1)}$ (5.21)
for each $\{I,J\}$.
Next we fix $\ell\in J$ and choose $j,j^{\prime}\in I$ ($j\neq j^{\prime}$). For each $\ell$ there are $\binom{k}{2}$ ways to choose such $j,j^{\prime}$. The probability to have $\langle j,\ell\rangle\in\mathfrak{E}$ and $\langle j^{\prime},\ell\rangle\in\mathfrak{E}$ for each triple $\{\ell,j,j^{\prime}\}$ is $p^{2}$. Thus we obtain
$\mathbb{P}\left(\exists\{j,j^{\prime}\},\langle j,\ell\rangle\in\mathfrak{E}\ {\rm and}\ \langle j^{\prime},\ell\rangle\in\mathfrak{E}\right)\leq\binom{k}{2}p^{2}$ (5.22)
for each $\ell$.
It can be seen from the above argument that
$\mathbb{P}\left(G_{n}\ {\rm satisfies}\ {\bf Condition\ 5.3}\right)$ (5.23)
$\leq$ $\sum_{k=2}^{\lfloor(n+1)/2\rfloor}\binom{n}{k}\binom{n-k}{k-1}\left\{\binom{k}{2}p^{2}\right\}^{k-1}(1-p)^{k(k-1)/2+k(n-2k+1)}$
$\leq$ $\sum_{k=2}^{\lfloor(n+1)/2\rfloor}\left(\frac{ne}{k}\right)^{k}\left(\frac{ne}{k-1}\right)^{k-1}\left\{\binom{k}{2}p^{2}\right\}^{k-1}{\rm exp}\left\{-pk\left(n-\frac{3}{2}k+\frac{1}{2}\right)\right\}$
$=$ $\frac{2}{nep^{2}}\sum_{k=2}^{\lfloor(n+1)/2\rfloor}\frac{1}{k}\left(\frac{(nep)^{2}}{2}\right)^{k}{\rm exp}\left\{-pk\left(n-\frac{3}{2}k+\frac{1}{2}\right)\right\}.$
When $2\leq k\leq 4$, one obtains
$\frac{2}{nep^{2}}\frac{1}{k}\left(\frac{(nep)^{2}}{2}\right)^{k}{\rm exp}\left\{-pk\left(n-\frac{3}{2}k+\frac{1}{2}\right)\right\}$ (5.24)
$=$ $O\left(\frac{\displaystyle(\log n)^{2(k-1)}}{\displaystyle n^{k-1}}\right)\rightarrow 0,\ \ \ n\rightarrow\infty.$
As we can see from $k\leq(n+1)/2$ that
${\rm exp}\left\{-p\left(n-\frac{3}{2}k+\frac{1}{2}\right)\right\}\leq{\rm exp}\left\{-\frac{p}{4}(n-1)\right\},$ (5.25)
the sum of the terms with $k\geq 5$ satisfies
$\varOmega$ $=$ $\frac{2}{nep^{2}}\sum_{k=5}^{\lfloor(n+1)/2\rfloor}\frac{1}{k}\left(\frac{(nep)^{2}}{2}\right)^{k}{\rm exp}\left\{-pk\left(n-\frac{3}{2}k+\frac{1}{2}\right)\right\}$ (5.26)
$\leq$ $\frac{1}{2np^{2}}\sum_{k=5}^{\lfloor(n+1)/2\rfloor}u^{k}\leq\frac{1}{2np^{2}}\frac{u^{5}}{1-u},$
where
$u=\frac{(nep)^{2}}{2}e^{-p(n-1)/4}.$ (5.27)
Now we take constants $c_{1},c_{2}\in\mathbb{R}$ complying with
$\frac{\log n+c_{1}}{n}\leq p(n)\leq\frac{\log n+c_{2}}{n}$ (5.28)
for sufficiently large $n$. It follows that
$\frac{1}{np^{2}}\leq\frac{n}{(\log n+c_{1})^{2}}$ (5.29)
and
$u\leq\frac{n^{2}e^{2}}{2}\left(\frac{\log n+c_{2}}{n}\right)^{2}{\rm exp}\left\{-\frac{\log n+c_{1}}{4}+\frac{p}{4}\right\}=\frac{e^{2+(p-c_{1})/4}}{2}\frac{(\log n+c_{2})^{2}}{\sqrt[4]{n}}.$ (5.30)
We can see from (5.26), (5.29) and (5.30) that
$\varOmega$ $\leq$ $\frac{n}{2(\log n+c_{1})^{2}}\frac{e^{10+5(p-c_{1})/4}}{32}\frac{(\log n+c_{2})^{10}}{n\sqrt[4]{n}}\left\{1-\frac{e^{2+(p-c_{1})/4}}{2}\frac{(\log n+c_{2})^{2}}{\sqrt[4]{n}}\right\}^{-1}$ (5.31)
$=$ $\frac{(\log n+c_{2})^{10}}{\sqrt[4]{n}(\log n+c_{1})^{2}}O(1)\rightarrow 0,\ \ \ n\rightarrow\infty.$
We are now in a position to asymptotically evaluate the probability for a symmetric random matrix $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$ to have $N$ distinct eigenvalues.
Theorem 5.5
Let $c\in\mathbb{R}$. Probability parameters $p,q$ depend on $N$ and satisfy
$p=p(N)=\frac{\log N+c}{N}+o\left(\frac{1}{N}\right),\ \ \ N\rightarrow\infty.$ (5.32)
$\lim_{N\rightarrow\infty}q(N)=q_{\infty}\in[0,1].$ (5.33)
Then we obtain
$\lim_{N\rightarrow\infty}\mathbb{P}\left({\rm the}\ {\rm matrix}\ (Y_{j\ell}Z_{j\ell})\ {\rm has}\ N\ {\rm distinct}\ {\rm eigenvalues}\right)=e^{-\mu}+\mu e^{-\mu},$ (5.34)
where $\mu=(1-q_{\infty})e^{-c}$.
Proof
It is sufficient to give the probability for the random bipartite graph $G=({\cal V}^{R},{\cal V}^{C},\mathfrak{E})$, $\mathfrak{E}=\{\langle j,\ell\rangle|Y_{j\ell}=1\}$ to satisfy Condition 4.1, when ${\cal E}=\mathfrak{E}$. Let us put $X=\#\{j\in{\cal I}_{N}|j\ {\rm is}\ {\rm an}\ {\rm isolated}\ {\rm point}\}$ and classify the cases according to the value of $X$.
(1) When $X\geq 2$, Condition 4.1 is not satisfied.
(2) When $X=0$, if Condition 5.3 is not satisfied, then Condition 4.1 is satisfied. It follows from Lemma 5.4 that
$\lim_{N\rightarrow\infty}\mathbb{P}({\bf Condition\ 5.3}\ {\rm is}\ {\rm satisfied})=0.$ (5.35)
This result and Lemma 5.1 lead to
$\lim_{N\rightarrow\infty}\mathbb{P}(X=0\ {\rm and}\ {\bf Condition\ 4.1}\ {\rm is}\ {\rm satisfied})=e^{-\mu}.$ (5.36)
(3) When $X=1$, suppose that $j\in{\cal I}_{N}$ is the isolated point. If the partial bipartite graph $G[{\cal I}_{N}\setminus\{j\}]$ has a perfect matching, then Condition 4.1 is satisfied. Let us consider each of the following two cases when $G[{\cal I}_{N}\setminus\{j\}]$ has no perfect matchings.
(a) First we consider the case when $G[{\cal I}_{N}\setminus\{j\}]$ has an isolated point $\ell\in{\cal I}_{N}\setminus\{j\}$, Since $\ell$ is not an isolated point of $G$, we obtain $\langle j,\ell\rangle\in\mathfrak{E}$. This is however contradictory to the fact that $j$ is an isolated point of $G$. Therefore it is impossible for $G[{\cal I}_{N}\setminus\{j\}]$ to have an isolated point.
(b) Next we consider the case when $G[{\cal I}_{N}\setminus\{j\}]$ satisfies Condition 5.3. Since $j\in{\cal I}_{N}$ is an isolated point of $G$, $G$ itself also satisfies Condition 5.3. One can see from Lemma 5.4 that the probability for $G$ to satisfy Condition 5.3 converges to zero in the limit $N\rightarrow\infty$.
The arguments on the cases (a) and (b) together with Lemma 5.1 yield
$\lim_{N\rightarrow\infty}\mathbb{P}(X=1\ {\rm and}\ {\bf Condition\ 4.1}\ {\rm is}\ {\rm satisfied})=\mu e^{-\mu}.$ (5.37)
Then we can conclude from the consideration of all the cases (1), (2) and (3) that
$\lim_{N\rightarrow\infty}\mathbb{P}({\bf Condition\ 4.1}\ {\rm is}\ {\rm satisfied})=e^{-\mu}+\mu e^{-\mu}.$ (5.38)
This is equal to the asymptotic probability for a symmetric random matrix $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$ to have $N$ distinct eigenvalues. ∎
6 Discussion
In this paper, the eigenvalue degeneracy probability is treated, when the probability measure of the random matrix entries is continuous except at the origin. As a result, we obtain a positive degeneracy probability for the zero eigenvalue of a sparse random matrix model. On the other hand, when the probability measure has discontinuous points other than the origin, similar results are not known. This limitation should be overcome by analyzing more general discontinuous models.
Acknowledgement
The author sincerely thanks Taro Nagao for valuable discussions, comments, and generous support, which greatly contributed to the clarity and presentation of this paper.